### v2

# Language Models can Self-Improve at State-Value Estimation for Better Search

Abstract
Collecting ground-truth rewards or human demonstrations for multi-step reasoning tasks is often prohibitively expensive and time consuming, especially in interactive domains like web tasks. To address this bottleneck, we present self-taught lookahead (STL), a self-supervised method that leverages state-transition dynamics to improve a value model capable of effectively guiding language model-controlled search without any labeled data. We find that moderately sized (8 billion parameters) open-weight value models improved with STL can match the performance of using a gpt-4o value model. Furthermore, we find that specialized value models learned with STL can be deployed with computationally lightweight search algorithms, achieving performance that matches that of more expensive tree search methods, while reducing costs by an order of magnitude.
1 Introduction
While large language models (LLMs) demonstrate strong reasoning capabilities by generating extended token sequences before answering wei2022chain ; guo2025deepseek ; muennighoff2025s1 , guiding inference with explicit tree search has the potential to further improve performance on tasks with a structured state space yao2024tree ; xie2024self . In this setting, a policy LLM proposes candidate actions, and a value LLM evaluates resulting states to steer the search toward promising trajectories. Figure 1 summarizes the assumptions different LLM-driven search methods make about information available during training and inference. Reward and Demo Learning methods yao2022webshop ; zhang2024rest_mcts ; jin2025search ; song2025r1 assume access to ground-truth reward signals or human demonstrations and optimize the model with reinforcement learning (RL) or imitation learning (IL). In contrast, Reward-Guided Inference strategies zhou2023lats ; shinn2024reflexion forego explicit reward during training; they only consult a reward signal at inference time, using it to guide procedures such as LLM-based Monte-Carlo Tree Search (MCTS). However, collecting ground-truth rewards or human demonstrations may not be possible in every environment, and can oftentimes be costly. For example, for web agent tasks, even small-scale data collection can cost thousands of dollars yao2022webshop . Reward and Demo Free methods yao2022react ; yao2024tree relax this assumption as they can operate without access to reward. However, these methods often rely on prompting an off-the-shelf LLM to serve as both the policy and value models during the search process, which constrains performance to the capabilities of the underlying LLM.
In this paper, we introduce Self-Taught Lookahead (STL), a Reward and Demo Free, self-supervised framework for interactive, multi-step reasoning tasks. Building on evidence that search quality is strongly influenced by the accuracy of the state-value estimator (chen2024search_useful, ; liu2024don, ), STL improves an LLM-based value function for better search performance. Unlike neural models traditionally used for state-value estimation in the learning to search literature silver2017masteringall , an LLM can leverage both conventional numerical values and natural language reasoning to estimate state values. STL exploits this feature by having an LLM value model learn to better assign values to states based on their expected future utility by constructing and learning from rationales that explicitly capture state transition dynamics. For instance, without understanding these transitions, it might not be clear whether a Close (X) button on a website interface exits the current view or the entire workflow in web tasks harley2019cancel . Learning better state-value estimates from state transition dynamics is especially well-suited for self-improvement in agentic tasks, where the environment directly provides transition outcomes. As a result, our approach requires neither ground truth rewards nor human demonstrations.
STL (Figure 2) begins by generating self-improvement data through a single step of lookahead within a tree search. Analogous to the Bellman update, this lookahead refines the estimated value of a state by leveraging information about potential future states. However, unlike classical reinforcement learning (RL) methods, which rely on explicit environment rewards, STL uses a large language model (LLM) to estimate state values. Specifically, during STL, a value LLM is fine-tuned to reason about the utility of a state by predicting the next best action, its resulting state, and a corresponding rationale for the value of that state. During training, the model is fine-tuned using rollouts of states and actions within the environment. At inference time, instead of taking a step of lookahead in the environment, the improved value model simulates a step of lookahead to provide more accurate value judgments.
By representing the lookahead process in natural language rather than regressing solely on value estimates, STL takes advantage of LLMs’ strong generalization to unseen tasks via learned textual reasoning nye2021show ; zelikman2022star . For instance, our results (§4) on web agent tasks from WebShop yao2022webshop demonstrate that tree search with an STL fine-tuned llama-3.1-8b-instruct value model improves performance by $39\%$ or more compared to the base llama model. Furthermore, STL matches the performance of search with a base gpt-4o value model and even achieves comparable results to Reward-Guided Inference methods such as LATS zhou2023lats on unseen tasks. We show that these results also hold for math puzzles and multi-hop question answering yao2024tree ; yang-etal-2018-hotpotqa . Finally, by enabling the use of a small open-source value model during inference, STL leads to significant cost reduction when agents are deployed. Through an efficiency analysis (§5), we find that STL costs 5$\times$ less than a similarly performing gpt-4o value model. As STL produces more accurate state-value estimates, the resulting value models can effectively guide search algorithms that expand fewer states, while maintaining strong performance.
Figure 1: The information accessible during learning and inference across common search settings, exemplified using web tasks. Our Self-Taught Lookahead method is Reward and Demo Free, yet is able to self-improve by learning from state transitions in the form of lookahead values and rationales.
Figure 2: Self-taught lookahead self-improves the value model by learning from state-transition dynamics. During the data generation phase (top left), tree search is used to discover diverse states. For every observed state $s$ encountered during the search, successor states are expanded using base policy $\pi_{\theta}$ and the current value model $V_{\phi_{k}}$, and a textual training example is formed using verbal representations of the next best action and successor state, as well as $V_{\phi_{k}}$’s outputted value reasoning ($r$) and numerical value ($v$) discounted by $\gamma$ (top middle). These examples are used to fine-tune $V_{\phi_{k+1}}$, which will be used in the next iteration of the algorithm (top right). Value models learned during STL can be used to evaluate unseen states encountered during search on unseen tasks by simulating a step of lookahead, including the next best action and the best successor state $\tilde{s}^{\prime}$ (bottom).
2 Background: Guiding Tree Search with Language Models
Within a state space $\mathcal{S}$, the goal of the tree search is to reach a desired state $s_{*}\in\mathcal{S}$ from an initial state $s_{0}\in S$, where $s_{*}$ is determined based on the natural language task $x$, with $x\in\mathcal{X}$, the set of all possible tasks. A state $s_{i}\in S$ might be a step in a reasoning chain or an intermediate webpage in a web navigation task. While the algorithmic details vary based on the tree search method e.g., breadth-first search (BFS) or MCTS, to adapt these methods to utilize language models (LMs), we simply need to define how new states (successors) are generated and evaluated.
Action generation.
Given a trajectory of $i+1$ states during the search process, candidate actions $a^{(j)}_{i}$ in the action space $\mathcal{A}$ are sampled using an LLM-based policy $\pi_{\theta}$:
$a^{(j)}_{i}\sim\pi_{\theta}(a_{i}|x,s_{0},\ldots,s_{i}),\forall j\in\{1,\ldots,B\}$ (1)
where $B$ is the branching factor or the number of specified candidate actions. These sampled actions constitute the set $A_{s_{i}}$. We denote the transition function as $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, so for an action $a_{i}$, $s_{i+1}=T(s_{i},a_{i})$.
State evaluation.
A value $v_{s_{i}|x}$ and a rationale for the value $r_{s_{i}|x}$ of a state $s_{i}$ is generated using a LLM-based value model $V_{\phi}:\mathcal{X}\times\mathcal{S}^{*}\rightarrow\mathbb{R}\times\mathcal{L}$, where $\mathcal{S}^{*}$ is the set of all finite sequences over the state space $S$ and $\mathcal{L}$ is the space of natural language sequences:
$(r_{s_{i}|x},v_{s_{i}|x})\sim V_{\phi}(x,s_{0},\ldots,s_{i})$ (2)
Note that because $V_{\phi}$ is an LLM, it is usually prompted in a chain-of-thought wei2022chain manner to first generate $r_{s_{i}|x}$ and then conditionally generate $v_{s_{i}|x}$ based on the rationale. For notational simplicity, will denote these two generated entities as $r_{s_{i}|x}\sim\text{Rat}(V_{\phi}(x,s_{0},\ldots,s_{i}))$ and $v_{s_{i}|x}\sim\text{Val}(V_{\phi}(x,s_{0},\ldots,s_{i}))$. While previous work zhou2023lats ; koh2024tree ; yao2024tree ; yu2024exact do generate rationales during state evaluation, usually they are generated purely to leverage the performance improvements of LLMs when asked to rationalize and are subsequently discarded. As described in §3, inspired by self-taught reasoning zelikman2022star , we explicitly use these rationales during self-improvement by fine-tuning on them along with the value to learn how to better estimate state values in a given domain.
3 Better State-Value Estimation with Self-Taught Lookahead
In this section, we present our proposed self-taught lookahead method. §3.1 and §3.2 describe the STL improvement procedure, while §3.3 explains how a STL-improved value model operates during inference. See Figure 2 and Algorithm 1 in Appendix A for an overview of the method.
3.1 Generating Rollouts
STL assumes a static policy model $\pi_{\theta}$ and only trains the value model through one or more iterations of self-improvement. We denote the value model initialized with a base LLM $V_{{\phi}_{0}}$ and the value model used in a subsequent iteration $k$ to assign values and generate rationales $V_{\phi_{k}}$.
An iteration $k$ of STL starts with a dataset $\mathcal{D}_{\texttt{rollout}_{k}}\subset\mathcal{X}$ of natural language tasks for the current iteration. For each $x_{i}\in\mathcal{D}_{\texttt{rollout}_{k}}$, we roll out the search tree using $\pi_{\theta}$ and $V_{\phi_{k}}$. Using tree search enables us to collect a diverse set of states so that the value model trained on these states’ values can better generalize to unseen states and tasks. We demonstrate this generalization in §4. When visiting state $s_{j}$ on the trajectory $\{s_{0},\ldots,s_{j}\}$ during tree search, we compute $s_{j}$’s lookahead value, $y_{s_{j}}$:
$y_{s_{j}}\leftarrow\gamma\max_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big{\}}$ (3)
where $\gamma$ is the discount factor. Since tasks are episodic, we set $\gamma=1$ bernstein2001reinforcement ; van2005survey ; leffler2007efficient . These lookahead values capture a better estimate of the true value of $s_{j}$ as they account for $s_{j}$’s successor states. In §6, we describe how generating and learning from these $y$’s is similar to fitted value iteration.
Action-outcome rationales.
However, alone, these lookahead values fail to reflect why a given state is valuable as they do not capture (1) which action yielded the best (highest value) successor state and (2) why the best successor state was assigned a high value by $V_{\phi}$.
To better capture the state transition dynamics, we also generate action-outcome rationales when visiting a state $s_{j}$. These rationales are of the form “{action} {outcome_state} {value_rationale}” where action is selected by the $\max$ operator in Equation 3, outcome_state is the state observed after taking the action in the environment and value_rationale is the rationale for the evaluation of this successor state generated by $V_{\phi_{k}}$. Fine-tuning on these rationales will enable a value model to predict the result of taking an action and incorporate this prediction (lookahead) into the current state’s value estimate. Formally, we can define these action-outcome rationales $o_{s_{j}}$:
$o_{s_{j}}\leftarrow a^{*}_{j}||s^{*}_{j+1}||\text{Rat}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},s^{*}_{j+1}))$ (4)
where |$\cdot$|$\cdot$ | | ⋅ denotes concatenation, $s^{*}_{j+1}\leftarrow T(s_{j},a^{*}_{j})$, and
$a^{*}_{j}\in\operatorname*{arg\,max}_{a\in A_{s_{j}}}\Big{\{}\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big{\}}\$ (5)
The training data set for iteration $k$ is thus a set of tuples: $\mathcal{D}_{k}=(s_{k},o_{s_{k}},y_{s_{k}})$. Depending on the task, it might be necessary to automatically filter out tuples that have malformed rationales or account for the same state seen multiple times in different iterations (see Appendices DE, and F).
3.2 Fine-Tuning the Value Model
We start training the new value model $V_{\phi_{k+1}}$ from the initial or base LLM value model $V_{\phi_{0}}$. We can then train using standard fine-tuning negative log-likelihood loss for the generation of both the action-outcome rationale and the lookahead value ($o_{s}||y_{s}$) of the state. We train the value model to generate the rationale before estimating the value. Automatically constructed text or formatting, as seen in Figure 2, is applied for easier learning.
3.3 Search after Self-Taught Lookahead
A value model resulting from iteration $k$ of STL ($V_{\phi_{k+1}}$) can directly replace a value model in any search algorithm, such as Greedy Search (§4.1 and §4.2) and BFS (§4.3). As shown in Figure 2 , $V_{\phi_{k+1}}$ simulates a step of lookahead for the state $s_{n}$ i.e. for $({r}_{s_{n}|x},{v}_{s_{n}|x})\sim V_{\phi_{k+1}}(x,s_{0},\ldots,s_{n})$,
${r}_{s_{n}|x}=\tilde{a}_{{n+1}}||\tilde{s}_{{n+1}}||\tilde{r}_{s_{n+1}|x}$ (6)
where $\tilde{a}_{{n+1}}||\tilde{s}_{{n+1}}$ is a simulated lookahead step and $\tilde{r}_{s_{n+1}|x}$ is its value rationale.
4 Experiments
We benchmark our proposed STL self-improvement approach on applied web agent tasks, multi-step question answering, and math puzzle tasks [^1] Our code is available at https://github.com/ethanm88/self-taught-lookahead..
4.1 Web Tasks
As mentioned in §1, it is particularly challenging and expensive to gather ground truth web task completion data zhang2024rest_mcts . To benchmark our STL method on web tasks, we utilize WebShop yao2022webshop , which consists of interactive web tasks involving searching for and purchasing an item that matches a short natural language specification. This benchmark is an ideal test bed to demonstrate the ability of our approach, as, unlike other web task datasets zhou2023webarena ; koh2024visualwebarena , ground truth reward is provided for all tasks, allowing a direct comparison between STL and methods in the Reward and Demo Learning and Reward-Guided Inference settings which use this reward.
[…]
Setting Method Policy Value Mini Test Set (50) Full Test Set (500)
Score $\uparrow$ SR $\uparrow$ Score $\uparrow$ SR $\uparrow$
[…]
Reflexion gpt-3.5-turbo —— 58.5 24.0 49.5 16.4
[…]
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 70.0 26.0 67.7 26.4
gpt-3.5-turbo r1-distill-llama-8b 68.4 24.0 66.3 24.6
gpt-3.5-turbo gpt-3.5-turbo 71.5 38.0*** 70.6*** 35.6***
gpt-3.5-turbo gpt-4o 72.9* 42.0*** 71.5*** 40.6***
gpt-4o llama-3.1-8b-instruct 71.6 28.0 67.2 25.8
gpt-4o r1-distill-llama-8b 71.6 32.0* 66.5 25.6
gpt-4o gpt-3.5-turbo 77.4*** 46.0*** 72.4*** 38.8***
gpt-4o gpt-4o 74.4** 46.0*** 71.4*** 40.8***
[…]
Human Expert —— —— —— 76.1 54.0 82.1 59.6
STL for web tasks.
Following the empirical advantages on agent tasks identified by yu2024exact , we generate training data with MCTS by performing a step of lookahead at each step during rollout. Note that we use the LLM value model value outputs as a proxy reward to guide UCT (Upper Confidence bounds applied to Trees) kocsis2006bandit selection like yu2024exact , instead of ground truth reward zhou2023lats . We perform STL with a gpt-3.5-turbo gpt-35 policy to be consistent with previous work zhou2023lats ; shinn2024reflexion as well as a gpt-4o GPT-4o policy and fine-tune a llama-3.1-8b-instruct dubey2024llama value function. STL is performed by rolling out 50 tasks from the WebShop training set, resulting in close to 1500 training examples, which we find is sufficient for significant performance improvement. Also, while we perform data generation during STL with MCTS, we evaluate the agent using the trained value models with greedy search, where the next action is greedily chosen based on the values of the policy’s proposed actions. Finally, we find that a single iteration of STL is sufficient to see significant improvement over using a base LLM-initialized value model, and that multiple iterations do not yield additional performance improvements due to difficulties in simulating more than one step ahead, given the complexity of the environment. This conclusion has been corroborated by other LLM agent works gu2024your ; chae2024web . However, in §4.3 we show that this multi-step simulation is possible for simpler tasks. See Appendix D for further details about data generation, training, and simulation.
Baselines.
Within Reward and Demo Learning approaches, we include IL and RL methods originally proposed in the WebShop work that train BERT devlin2019bert and BART lewis2019bart models on human demonstrations and ground truth reward yao2022webshop . We additionally compare to the current state-of-the-art approach AgentQ putta2024agent , which finetunes a larger xLAM-v0.1-r-46.7b policy on rolled out MCTS search trees using direct preference optimization rafailov2023direct . We also include Reward-Guided Inference methods such as Reflexion shinn2024reflexion and LATS zhou2023lats , which work by prompting closed-source LLMs within a framework, e.g., MCTS, that is guided by ground truth reward. Finally, in the Reward and Demo Free setting, we include greedy search and MCTS [^2] We use LLM value as a proxy reward to guide UCT, like in the data generation phase of STL baselines with a ReACT yao2022react prompted base LLM (llama-3.1-8b-instruct, gpt-3.5-turbo, gpt-4o, r1-distill-llama-8b guo2025deepseek ). All value models are prompted with few-shot examples and are asked to provide reasoning before a numerical value. For clear comparisons, a branching factor of 5 is used for greedy and MCTS-based methods like LATS, i.e., LATS is strictly more computationally expensive than greedy methods. We use the pass@$3$ chen2021passk for methods that do not have access to reward at inference time.
Results and discussion.
The WebShop average reward (Score) and success rate (SR) of all methods evaluated are presented in Table 1. We present results on both the full WebShop test set and on the mini test set of 50 tasks used by zhou2023lats , as we find running LATS and other MCTS methods on the entire test set is computationally expensive. Both of these sets are distinct from those seen during STL. Within the Reward and Demo Free setting, we find that STL matches the performance of using a gpt-4o value model and leads to a greater than 7% improvement in average reward and a 39% improvement in success rate relative to a base llama-3.1-8b-instruct value model, both of which are statistically significant improvements ($p<0.001$) using the paired bootstrap test berg2012empirical . Moreover, we find that STL even outperforms Reward-Guided Inference methods that have access to ground truth rewards. Lastly, we find no statistically significant difference between using a gpt-3.5-turbo or gpt-4o policy due to high action diversity in our setup (see Appendix D.2); thus, we evaluate search methods with a single policy (either gpt-3.5-turbo or gpt-4o) in subsequent experiments.
Table 2: Ablation study on the impact of fine-tuning with different combinations of information from lookahead namely lookahead values (lv), textual representation of the next best action and successor state (tr), and the value rationale for the successor state (r). The underlined results are from the base model before any fine-tuning.
Fine-tuning Data Setup Score $\uparrow$
llama-3.1-8b-instruct 70.0
+ lv 70.9
+ lv + tr 74.4
+ lv + tr + r (stl) 78.3
Reasoning Ablation.
In the Reward and Demo Free setting, we also perform ablations on the set of information from the step of lookahead used to fine-tune the value model during self-improvement. Specifically, we compare STL to variants that use only subsets of the information derived from lookahead. As mentioned in §3, this information includes (1) the lookahead value, (2) the textual representation of the next-best action and its successor state, and (3) the value rationale for the successor state. The results of this ablation in Table 2 demonstrate that regressing solely on lookahead values and further incorporating state transitions from lookahead does improve performance relative to the base model. However, learning also from the value rationale of the successor state, as done in STL, yields additional performance gains over these other settings. These results substantiate the claims made in §3 about the necessity of learning from action-outcome rationales, a key difference between STL and both classical RL gordon1999fitted_value_iteration and other LLM tree search works feng2023alphazero ; zhang2024rest_mcts which fine-tune an LLM value model on numerical values only.
4.2 Multi-Hop Question Answering
We also investigate the efficacy of STL on applied reasoning for retrieval-based question-answering. We specifically utilize the HotpotQA yang-etal-2018-hotpotqa benchmark, consisting of multi-hop question-answering tasks that require retrieving and reasoning over multiple Wikipedia entries. We use the same setup as in §4.1 to generate data by rolling out with MCTS, but instead roll out 500 tasks from the training dataset since actions (search terms) proposed by the policy lack diversity compared to web tasks.
Baselines.
In the Reward and Demo Learning setting, we evaluate on R1-Searcher song2025r1 , which performs RL on outcome-based answer correctness rewards on the multi-step retrieval QA task. As in §4.1, in the Reward-Guided Inference setting, we include Reflexion and LATS baselines, which have access to ground truth correctness during search. Among Reward and Demo Free methods, we compare STL to greedy search with closed-source models.
Results and discussion.
Table 3 presents the answer match rate of all evaluated methods. Following feng2024agile , we prompt gpt-4o to help evaluate answer correctness. The prompt used is in Appendix E.3. This fuzzy matching correctness is exposed as the reward for Reflexion and LATS, but not to the greedy methods in the Reward and Demo Free setting. We evaluate on a set of 500 unseen questions and also a smaller set of 50 examples due to the high cost of LATS. STL on llama-3.1-8b-instruct outperforms a gpt-4o value model and the R1-Searcher method while approaching the performance of Reward-Guided Inference methods, which can verify predicted answer correctness during inference. Of the Reward and Demo Free approaches, STL is the only method that is statistically significant compared to the llama-3.1-8b-instruct baseline ($p<0.001$).
[…]
Setting Method Policy Value Match Rate $\uparrow$
Test Set (50) Test Set (500)
[…]
Reflexion gpt-3.5-turbo 70.0 66
[…]
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 60.0 56.4
gpt-3.5-turbo gpt-3.5-turbo 62.0 56.0
gpt-3.5-turbo gpt-4o 68.0 57.6
[…]
4.3 Math Puzzles
Figure 3: BFS Game-of-24 performance on tasks seen and unseen during STL.
Finally, we also study the performance of STL on the Game-of-24 task yao2024tree , where the goal is to construct a mathematical expression with 4 provided integers to obtain 24. The task was originally for Reward and Demo Free methods, so it serves as a good benchmark for STL.
STL for Game-of-24.
For this task, we generate data using breadth-first search (BFS) rather than MCTS to be consistent with the original Tree-of-Thoughts yao2024tree approach. We use a gpt-4o policy and a llama-3.1-8b-instruct base value function in the first iteration. As described in §3, we replace this base value function with a trained model in each subsequent iteration. STL is run for four iterations of 25 puzzles. For this task, we do not have explicit environment observations, but instead use the policy’s arithmetic to combine two numbers as a pseudo-observation. See Appendix F for further details.
Baselines.
We compare the performance of value models learned via STL with Reward and Demo Free BFS baselines that use the same gpt-4o policy. Specifically, we experiment with a gpt-4o value model and an algorithmic oracle evaluator chen2024search_useful . This oracle runs a recursive algorithm to verify whether the current state (set of numbers) can be combined to reach $24$. Search performance with this oracle is an upper bound on the performance improvement possible from improving the value function under a static policy.
Results and discussion.
Figure 3 shows the performance of evaluated methods on a set of 50 tasks seen during STL and a set of 50 more challenging (determined by lower human solve percentages), unseen tasks. On both sets, STL matches or outperforms a gpt-4o value model. However, STL’s performance on seen tasks monotonically increases for the first three iterations, while its performance on unseen tasks decreases before increasing. This phenomenon may be due to the base value model generalizing better than self-improved models in earlier iterations, during which the model has not received signal from more accurate LLM values of terminal states during lookahead.
Figure 4: Compute and environmental efficiency during evaluation on WebShop with a gpt-3.5-turbo policy (left). Compute efficiency is measured in total (prompt and completion) tokens. Environmental efficiency is measured by the number of states expanded (webpages visited). The distribution of tokens (closed vs. open source models) used during search is also shown (right). Value models are specified in parentheses.
5 Efficiency Analysis
In this section, we compare the efficiency of STL with prior methods on WebShop. We study efficiency tradeoffs from two perspectives (1) model costs and (2) environment usage. Additionally, in §5.3 we explore how performance changes when scaling the size of the value model trained during STL.
5.1 Compute and Cost Efficiency
Keeping compute requirements and costs low is critical, especially for agents automating routine, repetitive tasks. Since STL can be used to improve an open-source LLM like llama-3.1-8b-instruct, we can transfer computation from more expensive closed-source models like gpt-4o to open-source alternatives while maintaining performance. Figure 4 (right) demonstrates this transfer, as STL uses more than 50% fewer closed-source tokens than greedy search with a gpt-3.5-turbo or gpt-4o value model. We also compute the monetary cost of different methods, accounting for both inference costs and, for STL, those incurred during fine-tuning. See Appendix G for details about the cost calculation. We plot these costs against the average WebShop reward in Figure 5 (left), and find that STL is Pareto optimal, 23×23$\times$23 × cheaper than MCTS methods like LATS, and 5×5$\times$5 × cheaper than performing greedy search with a gpt-4o value model.
Figure 5: Tradeoff between performance and efficiency on WebShop with a gpt-3.5-turbo policy. Pareto frontiers of existing methods and baselines are shown, illustrating the optimality of STL when considering the tradeoff between cost and average reward (left) and between environmental usage and average reward (right). Fine-tuning costs are included in the total STL cost.
5.2 Environmental Usage
It is often crucial for an agent taking actions in physical or digital environments to be environmentally efficient or conservative in the number of states it visits while performing a task. In the case of digital web agents, taking many steps per task through exhaustive tree search may put an unnecessary burden on web servers, especially as agents are deployed at scale. Additionally, allowing web agents to search widely when equipped with personal information or the ability to make purchases may lead to unintended privacy disclosures or financial loss, respectively. In some environments, taking many actions may also lead to unreasonable task completion times. Figure 4 (left) and Figure 5 (right) present the environmental efficiency measured by the count of expanded states or visited sites in the WebShop environment. Considering WebShop score, STL is Pareto optimal and requires expanding half as many states as MCTS-based methods like LATS. Moreover, unlike LATS, STL does not require irreversible actions (actually clicking Buy Now on a product page) required to obtain reward.
5.3 STL Scaling Trends
Figure 6: STL scaling trends on WebShop for llama-3 and qwen-2.5 families with a gpt-3.5-turbo policy.
As 8B STL models can match the performance of a gpt-4o value model, is it possible to use even smaller models for STL while maintaining good performance? STL requires models to (1) provide generally consistent values out-of-the-box so that it is possible to compare successor states during data generation and (2) learn to generalize to unseen tasks and states, both of which may be challenging for smaller models. We explore this STL scaling trend on WebShop with $\leq$ 8 billion parameter models in the llama 3 family [^3] As the llama-3.1 family lacks smaller models, we use 1B and 3B models from the llama-3.2 family. dubey2024llama and $\leq$ 7 billion parameter models in the qwen-2.5-instruct family yang2024qwen2 . The results presented in Figure 6 demonstrate that while performance does generally decrease with fewer parameters, the smaller 3B parameter models in both families do approach gpt-4o performance. This result suggests that using smaller models for STL is feasible, making large-scale agent deployment to new domains more practical.
6 Related Work
Classical RL.
STL is loosely inspired by fitted value iteration (FVI) gordon1999fitted_value_iteration , which generalized value iteration bellman1957dynamicprog to the tabular setting. In an iteration of FVI, target values are computed using the Bellman update and used to train a new value model from the previous model checkpoint using least squares regression. The iterated values in FVI are computed similarly to the lookahead values $y_{s_{k}}$ in §3.1, but with STL, no ground truth reward is assumed, the value model is non-Markovian, and actions are deterministic. Instead of learning directly from iterated values, with STL, they are concatenated with action-outcome rationales, and together, these sequences are used to fine-tune the LLM value model from scratch at each iteration rather than from the previous model checkpoint as in FVI.
LLM self-improvement.
A variety of previous work has shown that LLMs can self-improve with iterative prompting techniques huang2022large ; weng2022large ; madaan2024self and have applied these methods to various domains, from agents kim2024language to privacy protection chen2023can . A separate line of work focuses on bootstrapping a small training dataset through a self-training process to improve either the reasoning policy model gulcehre2023rest ; singh2023rest_em ; jung2023impossible or the verification or reward model hosseini2024v using synthetically generated data. While most self-training approaches utilize outcome-based reward models, other work aksitov2023rest_react ; zhang2024rest_mcts derive process-based rewards like STL to evaluate each step in the reasoning chain.
Training reasoning agents.
The majority of prior work on training reasoning agents focuses on performing SFT on human-annotated trajectories yao2022webshop ; lai2024autowebglm , synthetically generated trajectories chen2023fireact ; furuta2023multimodal ; zhang2024xlam ; liu2024apigen ; patel2024large ; murty2024nnetscape ; ou2024synatra ; su2025learn , or a combination of both zeng2023agenttuning ; zhang2024agentohana . Other work has trained agents from tree search generated data, but these methods usually require ground truth reward gandhi2024stream ; putta2024agent ; zhang2024rest_mcts . While prior work has explored self-improving reasoning agents, these approaches fail to generalize beyond the instructions encountered during self-improvement patel2024large or require fine-tuning frontier models like gpt-4o to achieve generalization yu2024exact .
7 Conclusion
We propose STL as an efficient method to improve the value model employed during search. This efficiency primarily stems from STL’s design for information-scarce settings, where models learn from state-transition dynamics. Additionally, because STL enables self-improvement on small models deployable with less exhaustive search, it yields significant reductions in both computational cost and environmental usage. Therefore, the STL framework could help enable the more realistic learning and deployment of agent systems.
Acknowledgments
We would like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This research is supported in part by the NSF under grant number IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Self-Taught Lookahead Algorithm
The STL algorithm is presented in full in Algorithm 1. For information about the task_specific_filter, see Appendix D.3 and Appendix F.4.
Algorithm 1 Self-Taught Lookahead
0:  Set of tasks $\mathcal{D}_{\texttt{rollout}}$, Base LLM $M$, num iterations $n$, num tasks per iteration $m$
  $\pi_{\theta}\leftarrow\texttt{initialize{\textunderscore}policy{\textunderscore}model}(M)$
  $V_{\phi_{0}}\leftarrow\texttt{initialize{\textunderscore}value{\textunderscore}model}(M)$
  $\mathcal{D}_{\texttt{Train}_{0}}\leftarrow\{\}$
  for $k\leftarrow 1$ to $n$ do
     $\mathcal{D}_{\texttt{rollout}_{k}}\leftarrow\mathcal{D}_{\texttt{rollout}}[m\cdot(k-1):m\cdot k]$Select tasks for iteration $k$}
     for $x\in\mathcal{D}_{\texttt{rollout}_{k}}$ do
        $T\leftarrow\texttt{rollout\_tree}(\pi_{\theta},V_{\phi_{k-1}},x)$Generate rollout search tree for task $x$}
        $y\leftarrow\texttt{calculate\_lookahead\_values}(T)$
        $o\leftarrow\texttt{calculate\_action\_outcomes\_rationales}(T)$
        $y_{\texttt{filtered}},o_{\texttt{filtered}}\leftarrow\texttt{task\_specific\_filter}(y,o)$If applicable, apply task-specific filtering}
        $\mathcal{D}_{\texttt{Train}_{k}}\leftarrow\texttt{add\_new\_data}(\mathcal{D}_{\texttt{Train}_{k}},(y_{\texttt{filtered}},o_{\texttt{filtered}}))$
     end for
     $V_{\phi_{k}}\leftarrow\texttt{fine\textunderscore tune}(V_{\phi_{0}},\mathcal{D}_{\texttt{Train}_{k}})$Finetune from base model}
  end for
Appendix B Broader Impacts
Our STL method enables LLMs to self-improve at search by capturing the mechanics of traditional RL algorithms in natural language. While we leverage this technique for better state-value estimation, it can likely be applied to other problems. Additionally, we show in §5, that STL requires significantly less compute overhead than similarly performing methods, thus reducing the energy consumption required for search during inference, helping to improve the sustainability of agent deployment.
As with all self-improvement work [43, 61], having models self-improve without human supervision may enable agents to learn how to take actions towards task completion that are not well-aligned with human values or preferences. To prevent misuse, we only release code to reproduce experiments in this paper and not to improve general-purpose agents for which these risks might be more prevalent. While investigating these harms is out of the scope of this work, we encourage future research in this area.
Appendix C Limitations
During self-improvement, STL does require task specifications like “I am looking for a queen-sized bed that is black …’ (see Figure 2) for web tasks. However, this assumption is well-founded as prior self-improvement work, such as that on LLM alignment [62, 54], also assumes tasks, or in their case, user prompts, are provided to initialize the self-improvement process. Additionally, as mentioned throughout the paper, this setting is much more reasonable than Reward and Demo Learning and Reward-Guided Inference settings, which require ground truth reward and/or human demonstrations along with task specifications.
Additionally, due to compute constraints, our experiments utilize models up to 8 billion parameters for STL. However, our results in §5 demonstrate a scaling trend in performance with respect to model size, indicating that using larger models may yield more performant agents. Given these positive results, we leave applying STL to larger models to teams with larger resource budgets.
Appendix D Self-Taught Lookahead on WebShop
In this section, we outline information about WebShop [58] and the implementation details of running STL on the benchmark.
D.1 WebShop Task
There are two main types of actions in the WebShop task:
  • search[query]:
    Search actions allow the user to search for a particular item with a natural language query, e.g., search[easy to use medium color face kit less than 40 dollars]. This action can only be taken on the search page, which is also the initial / home page of the WebShop interface.
  • click[button]: Click actions are discrete actions but can take many forms, which we enumerate below:
    • click[product]: to select a relevant product from the search results e.g. click[B09B6SH764] where B09B6SH764 is a product code.
    • click[attribute]: to toggle on an attribute or option on the product page of an item, e.g. click[small]
    • click[Buy Now]: to buy the selected item - this is a terminal action that yields the ground truth reward. This action is not allowed to be taken in STL search but is allowed in other search, RL, and prompting methods [58, 46, 68].
    • Other navigation buttons: other navigation buttons include click[Back to Search], click[<Prev], click[Next>], click[Description], click[Features]. To simplify trajectories, we generally restrict the ability for models to take these actions in all settings following [58].
WebShop provides a textual representation of webpages in simple mode. An example of this representation for search results is shown in Figure 7.
WebShop Textual Representation [Back to Search] Page 1 (Total results: 50) [Next >] [B0972Q1T8T] Cosycost USB Microphone, Condenser Computer PC Gaming Microphone for PS4/5 Laptop Windows Mac OS Android Phone, Noise Cancelling Instant Mute, Studio Mic for Voice, Music Recording, Podcasting, Streaming \$32.99 [B09N3M6H2Z] Wired Stereo Headset Noise Cancelling Microphone with in-line Controls/Volume Controller, All-Day Comfort Design, Works for Playstation, Nintendo Switch, PC with USB Connection (HS-HP101UNCBK) \$199.99 [B072L2D6LY] Andrea Communications NC-255VM USB On-Ear Stereo USB Computer Headset with Noise-Canceling Microphone, in-Line Volume/Mute Controls, and Plug \$34.59 [B071H84LTJ] Andrea Communications NC-455VM USB Over-Ear Circumaural Stereo USB Computer Headset with Noise-Canceling Microphone, in-Line Volume/Mute Controls, and Plug \$49.24 [B08GLJSWJ9] Jiade USB Headset with Noise Canceling Microphone for CallCenter Skype Chat, Computer Phone Headset Voice Recognition Speech Dictation, PC Headphone with Mic Mute Volume Control Binaural Golden \$9.99
Figure 7: Example of simple mode textual representation of the state with the WebShop benchmark.
D.2 Prompts
The prompt used to generate actions in WebShop is presented in Figure 8. Notice that we do not use think actions part of the classical ReACT framework [60] like [58] or [68] because evaluating the value of these actions is difficult as they have no observation. Instead, we prompt the policy model to provide a rationale while generating possible actions. Also, note that we prompt the policy multiple times, adding to the list of actions that are not allowed and removing from the list of actions that are allowed. This change to a “selection” policy enables action diversity, which we find is otherwise low even with prompting the policy at high temperature. This change also likely explains why there is no statistically significant difference between using a gpt-3.5-turbo and gpt-4o policy in §4.1.
Likewise, the prompt used to evaluate states is presented in Figure 9. Note that this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. We note that the Likert scale used was crucial to obtaining consistent value outputs on which we could perform STL. For all value estimates (base model or fine-tuned), we prompt the value model 5 times and use the average score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 5 scores.
WebShop Generation Prompt You are a web agent, select the best next action for the search to fulfill the task. Example tasks are shown below. Provide a rationale for your selection BEFORE you provide the action. NOTE: You can only select actions that are provided in the Possible Actions list. You MAY NOT select actions in the Not Allowed list. NOTE: You must output BOTH a rationale and an action. NOTE: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page. Example Tasks: {few shot examples} —————————————————————————————————– New Task: {task} Actions Not Allowed: {not_allowed_actions} Possible Next Actions (REMINDER: You can only select actions from this list.): {possible_actions} REMINDER: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page.
Figure 8: Generation prompt for WebShop policy.
WebShop Value Estimation Prompt Given an item to purchase and a trajectory that aims to buy an item that exactly matches the specification, analyze how well the last action and observation align with the task. Provide a reflection that concludes with. "Thus the correctness score is s", where s is either 1, 2, 4, 6, 8, or 10. Use the following scale for scoring: 1: The last action and observed state is entirely irrelevant to the task or captures a purchase of an item that is completely unrelated to the specifications. 2: The last action and observed state captures a step with a low likelihood of leading to purchasing the correct item. 4: The last action and observed state captures a step with a moderate likelihood of leading to purchasing the correct item. 6: The last action and observed state captures a step with a high likelihood of leading to purchasing the correct item. 8: The last action and observed state captures a step with a very high likelihood of leading to purchasing the correct item. 10: The last action and observed state captures a step that will certainly lead to purchasing the correct item. Keep reflections short (<100 words). Follow the format of the rationale from the below example task. NOTE: the observation from clicking on the item will be the item’s product detail page. For instance, click[B078GWRC1J] will show the product detail page for the item with code B078GWRC1J which will include the item’s name (e.g. Bright Citrus Deodorant by Earth Mama), price (\$10.99), and other relevant details as well as options. NOTE: Assume none of the attributes on the product page are selected only provide the reflection for the last action. Example Tasks: {few shot examples} —————————————————————————————————– New Task: Respond with the reflection for the last observation of the new task ONLY. As a reminder the last action and observation is as follows: {last_action} Your response should start with "Reflection:" and end with "Thus the correctness score is …".
Figure 9: Value Estimation prompt for WebShop. This prompt was only used to prompt base models.
D.3 Implementing STL
With web tasks, the position of an action in a trajectory may influence its value. For instance, selecting a certain item I from search results early in the trajectory should have a higher value than selecting I a second time in the same trajectory. To account for this difference, we train value models at each position (depth) in the trajectory. Specifically, we limit trajectories to five steps and train four value models depths 1 to 4, only allowing a terminating Buy action on the final step.
We also filter out malformed rationales from the training data. Specifically, we remove rationales that do not provide the proper format, e.g., it does not exactly contain scaffolding like “Thus the correctness score is”.
Additionally, we generate lookahead rollouts using 5 iterations of MCTS, which, in practice, we find is sufficient for collecting diverse training data.
D.4 Difficulty in Performing STL for multiple iterations.
Empirically, we find that STL after a second iteration on WebShop has a lower performance (average reward of 68.6, and success rate of 26.0) than after a single iteration. From a manual inspection of the lookahead and rationales generated, we notice that the second step simulated by the value model often does not match the true environment.
Appendix E Self-Taught Lookahead on HotpotQA
E.1 HotpotQA Task
The HotpotQA [57] is a multi-hop question answering benchmark where correct answers require reasoning over multiple Wikipedia entries. There are three possible actions at each step:
  • search[entry]:
    Search provides the first five sentences of the corresponding Wikipedia entry if it exists, or provides five alternative existing entries.
  • lookup[string]:
    Lookup returns the next sentence in the entry containing the specified string.
  • finish[string]:
    Finish signifies the completion of the reasoning process, where the answer is specified with the provided string.
E.2 Prompts
The prompt used to generate actions for the HotpotQA task is presented in Figure 12. Likewise, the prompt used to evaluate states is presented in Figure 11. Many of the prompting details, such as the use of a disallowed action list, are similar to the WebShop task. However, one difference is that if five unique actions are not found due to a search term not being an entry, we add the top-5 similar terms returned by the retrieval model as possible actions.
E.3 Evaluation
We use the same gpt-4o evaluation prompt as [13] which is provided in Figure 11. Note that a match is computed when either there is an exact match or a match indicated by the gpt-4o evaluator in order to prevent any mistakes by the evaluator, such as indicating a non-match even though the two answers were an exact string match.
Additionally, unlike WebShop, we limit trajectories to a depth of four, and explicitly prompt models to provide a final answer on the fourth step in the trajectory.
E.4 Implementing STL
The implementation details of STL for HotpotQA are similar to WebShop in the training of multiple models at different depths, the filtering out of malformed rationales, and the use of 5 iterations of MCTS during data generation. However, one difference is that we feed the possible next actions to the value model so that it can provide coherent lookahead simulations and evaluations.
HotpotQA Generation Prompt Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. After each observation, provide the next Thought and next Action. NOTE: You MAY NOT select actions in the Actions Not Allowed list. You have to change the wording of the query or lookup somehow. NOTE: Keep search queries and lookups short and concise since longer queries will not return any result. For instance, instead of searching Search[eastern sector of the Colorado orogeny], search Search[Colorado orogeny] and then Lookup[eastern sector]. Here are some examples: {few shot examples} —————————————————————————————————- New Task: {task} Actions Not Allowed: {not_allowed_actions} Thought:
Figure 10: Generation prompt for the HotpotQA policy.
HotpotQA Value Estimation Prompt Given a question and a trajectory to answer the question, analyze how well the LAST ACTION in the trajectory contributes to finding the answer. Consider ONLY the last action. The trajectories are labeled by pairs of thoughts that can reason about the current situation and actions that can be of three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. Provide a reflection that concludes with “Thus the correctness score is s”, where s is either 1, 3, 5, 7, or 10. Use the following scale for scoring: 1: The action is completely irrelevant to answering the question or there is no relevant search result (“Could not find” is in the observation). 3: The action’s observation provides information only at the background level to answering the question. 5: The action’s observation provides information that makes a small step towards answering the question. 7: The action’s observation provides information that makes a key step towards answering the question. 10: The action’s observation provides the final piece of information needed to answer the question. Reminder: If you see “Could not find” in the observation, the correctness score is 1. Keep reflections short (< 100 words). Follow the following examples: {few shot examples} —————————————————————————————————- {input}
Figure 11: Value estimation prompt for HotpotQA. This prompt was only used to prompt base models.
HotpotQA Evaluation Prompt Based on the provided question and reference answer, please determine if the response is correct or incorrect. Begin by articulating your rationale, and conclude with a single word judgment: ‘Yes’ for correct or ‘No’ for incorrect. question: {question} reference answer: {reference} response: {response}
Figure 12: Evalaution prompt for the HotpotQA predicted answers.
Game-of-24 Generation Prompt Use numbers and basic arithmetic operations (+ - * /) to obtain 24. In each step, you are only allowed to choose two of the remaining numbers to obtain a new number. Follow the example format exactly. {few shot examples} —————————————————————————————————- {input}
Figure 13: Generation prompt for the Game-of-24 policy.
Game-of-24 Value Estimation Prompt Evaluate if the given numbers can reach 24 (sure/likely/impossible) Follow the example format exactly. Only evaluate the last example. {few shot examples} —————————————————————————————————- {input}
Figure 14: Value estimation prompt for Game-of-24. This prompt was only used to prompt base models.
Appendix F Self-Taught Lookahead on Game-of-24
F.1 Game-of-24 Task
As introduced by [59], the Game-of-24 is a mathematical reasoning task that involves combining four numbers e.g. “2 3 4 5” together with mathematical operations i.e. $+$, --, ///, $\times$ in order to obtain 24. An action in this task consists of simply applying a mathematical operation to combine two numbers e.g. $2+3=5$, the resulting state from the operation is the set of remaining numbers e.g “5 4 5”.
F.2 Prompts
The prompt used to generate actions for the Game-of-24 task is presented in Figure 13. Likewise, the prompt used to evaluate states is presented in Figure 14. Like WebShop, this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. However, unlike WebShop, values are not real numbers between 1 and 10, but rather 0.001, 1, and 20, corresponding to the labels of impossible, likely, and sure that the remaining numbers can be combined to reach 24. Note that these values were used by the original Tree-of-Thoughts paper [59], but are ad-hoc and are used purely as labels. For all value estimates (base model or fine-tuned), we prompt the value model 3 times and use the median score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 3 scores.
F.3 Evaluation
For all tested BFS methods, we use the same setup as the Tree-of-Thoughts paper, i.e., we select 5 of the best actions (set the branching factor to 5) at each of the 4 steps (two numbers are combined during each step).
F.4 Implementing STL
Since the state space is quite limited, we combine training examples from the previous $k-1$ iterations with current examples to train the value model in the $k$th iteration. If the same state is encountered multiple times in different iterations, we defer to the value judgment from the latest iteration.
Unlike with WebShop and HotpotQA, we do not train value models at each depth, due to the small state space. Additionally, we also use a branching factor of 5 during data generation.
Appendix G Costs
Here we detail how we computed costs in the efficiency analysis in §5.
G.1 Inference Costs
To compute inference cost, we use OpenAI’s 444openai.com/api/pricing for OpenAI models like gpt-3.5-turbo and gpt-4o and Groq’s 555groq.com/pricing pricing tables for closed source llama models. Note that we only use Groq for a pricing estimate, as inference and training were run on in-house GPUs. As costs may change over time, we provide the pricing figures that we used for our experiments in Table 4.
Table 4: Inference pricing used for cost analysis.
Prompt Tokens (\$ / 1000 tokens) Completion Tokens (\$ / 1000 tokens)
gpt-3.5-turbo 0.0005 0.0015
gpt-4o 0.0025 0.01
llama-3.1-8b-instruct 0.00005 0.00008
G.2 Fine-tuning Costs
We also account for the costs incurred from fine-tuning llama-3.1-8b-instruct during STL. Fine-tuning with a single A40 GPU takes 4.5 hours for the WebShop task. While it is difficult to estimate costs on our own in-house GPUs, we can estimate the cost by using VastAI’s figure of \$0.39 per A40 GPU hour  [^6] https://vast.ai/pricing/gpu/A40 when these experiments were run in April, 2025. Thus, the total training run cost roughly \$1.76. We also note that as the number of tasks seen at test time scales, the fine-tuning costs become increasingly negligible.
Appendix H Model Fine-tuning and Serving
Table 5: Hyperparameters during STL training.
warmup-steps learning-rate weight-decay per-device-batch size lora-r lora-alpha
STL $10$ $2e^{-4}$ $0.01$ $8$ $16$ $16$
Table 6: Effect of $\gamma$ on Performance.
$\gamma$ $0.75$ $0.80$ $0.85$ $0.90$ $0.95$ $0.99$ $1.0$
WebShop Success Rate $32.0$ $24.0$ $24.0$ $24.0$ $28.0$ $42.0$ $46.0$
Fine-tuning the value model for STL is carried out on a single NVIDIA A40 GPU. We use LoRA finetuning [23] and use models provided by unsloth777unsloth.ai. The hyperparameters used are in Table 5. We fine-tuned Game-of-24 value models for 10 epochs and WebShop value models for 20 epochs due to the differences in difficulty for models to learn the format of the action and state representations.
Additionally, we serve base and fine-tuned models using vLLM 888docs.vllm.ai/en/latest/ [31] for efficient value estimation of new states during search.
Finally, we perform a small analysis on the effect of the discount factor $\gamma$ on search performance. The results of the analysis with a gpt-3.5-turbo policy on WebShop are presented in Table 6. These results confirm the choice of $\gamma=1.0$ from prior work cited in §3.
Appendix I Significance Testing
In §4.1 and §4.2, we use the paired bootstrap test to test the statistical significance of our experimental results. Following [3], we set $b=10^{6}$. For WebShop, we run the significance test twice: once for score (average reward) and a separate time for success rate.