### v3

# Language Models can Self-Improve at State-Value Estimation for Better Search

Abstract
Collecting ground-truth rewards or human demonstrations for multi-step reasoning tasks is often prohibitively expensive, particularly in interactive domains such as web tasks. We introduce Self-Taught Lookahead (STL), a reward-free framework that improves language model–based value functions by reasoning explicitly about state transitions. STL can be viewed as a chain-of-thought analogue of the value iteration algorithm: instead of regressing directly on numeric values, a value LLM is trained to simulate a step of lookahead in natural language—predicting the next action, resulting state, and rationale for its value, thereby refining value estimates without any labeled data. This self-supervised procedure yields more accurate state-value predictions, which in turn enable lightweight search algorithms to expand fewer states while maintaining strong performance. Empirically, STL-trained value models built on moderately sized (8B parameter) open-weight LLMs boost web agent success rates by 39%, achieving comparable performance with proprietary models. STL also generalizes to multi-hop QA and math puzzles. We find that STL enables small open-source models to guide efficient search, reducing inference costs by integrating explicit reasoning with value learning.
1 Introduction
While large language models (LLMs) demonstrate strong reasoning capabilities by generating extended token sequences before answering wei2022chain ; guo2025deepseek ; muennighoff2025s1 , guiding inference with explicit tree search has the potential to further improve performance on tasks with a structured state space yao2024tree ; xie2024self . In this setting, a policy LLM proposes candidate actions, and a value LLM evaluates resulting states to steer the search toward promising trajectories. Figure 1 summarizes the assumptions different LLM-driven search methods make about information available during training and inference. Reward and Demo Learning methods yao2022webshop ; zhang2024rest_mcts ; jin2025search ; song2025r1 assume access to ground-truth reward signals or human demonstrations and optimize the model with reinforcement learning (RL) or imitation learning (IL). In contrast, Reward-Guided Inference strategies zhou2023lats ; shinn2024reflexion forego explicit reward during training; they only consult a reward signal at inference time, using it to guide procedures such as LLM-based Monte-Carlo Tree Search (MCTS). However, collecting ground-truth rewards or human demonstrations may not be possible in every environment, and can oftentimes be costly. For example, for web agent tasks, even small-scale data collection can cost thousands of dollars yao2022webshop . Reward and Demo Free methods yao2022react ; yao2024tree relax this assumption as they can operate without access to reward. However, these methods often rely on prompting an off-the-shelf LLM to serve as both the policy and value models during the search process, which constrains performance compared to models specifically tuned for agentic tasks wei2025browsecomp .
In this paper, we introduce Self-Taught Lookahead (STL), a Reward and Demo Free, self-supervised framework for interactive, multi-step reasoning tasks. Building on evidence that search quality is strongly influenced by the accuracy of the state-value estimator (chen2024search_useful, ; liu2024don, ), STL improves an LLM-based value function for better search performance. Unlike neural models traditionally used for state-value estimation in the learning to search literature silver2017masteringall , an LLM can leverage both conventional numerical values and natural language reasoning to estimate state values. STL exploits this feature by having an LLM value model learn to better assign values to states based on their expected future utility by constructing and learning from rationales that explicitly capture state transition dynamics. For instance, without understanding these transitions, it might not be clear whether a Close (X) button on a website interface exits the current view or the entire workflow in web tasks harley2019cancel . Learning better state-value estimates from state transition dynamics is especially well-suited for self-improvement in agentic tasks, where the environment directly provides transition outcomes. As a result, our approach requires neither ground truth rewards nor human demonstrations.
STL (Figure 2) begins by generating self-improvement data through a single step of lookahead within a tree search. Analogous to the Bellman update, this lookahead refines the estimated value of a state by leveraging information about potential future states. However, unlike classical reinforcement learning (RL) methods, which rely on explicit environment rewards, STL uses a large language model (LLM) to estimate state values. Specifically, during STL, a value LLM is fine-tuned to reason about the utility of a state by predicting the next best action, its resulting state, and a corresponding rationale for the value of that state. During training, the model is fine-tuned using rollouts of states and actions within the environment. At inference time, instead of taking a step of lookahead in the environment, the improved value model simulates a step of lookahead to provide more accurate value judgments.
[…]
Figure 1: The information accessible during learning and inference across common search settings, exemplified using web tasks. Our Self-Taught Lookahead method is Reward and Demo Free, yet is able to self-improve by learning from state transitions in the form of lookahead values and rationales.
Figure 2: Self-taught lookahead self-improves the value model by learning from state-transition dynamics. During the data generation phase (top left), tree search is used to discover diverse states. For every observed state $s$ encountered during the search, successor states are expanded using base policy $\pi_{\theta}$ and the current value model $V_{\phi_{k}}$, and a textual training example is formed using verbal representations of the next best action and successor state, as well as $V_{\phi_{k}}$’s outputted value reasoning ($r$) and numerical value ($v$) discounted by $\gamma$ (top middle). These examples are used to fine-tune $V_{\phi_{k+1}}$, which will be used in the next iteration of the algorithm (top right). Value models learned during STL can be used to evaluate unseen states encountered during search on unseen tasks by simulating a step of lookahead, including the next best action and the best successor state $\tilde{s}^{\prime}$ (bottom).
2 Background: Guiding Tree Search with Language Models
Within a state space $\mathcal{S}$, the goal of the tree search is to reach a desired state $s_{*}\in\mathcal{S}$ from an initial state $s_{0}\in S$, where $s_{*}$ is determined based on the natural language task $x$, with $x\in\mathcal{X}$, the set of all possible tasks. A state $s_{i}\in S$ might be a step in a reasoning chain or an intermediate webpage in a web navigation task. While the algorithmic details vary based on the tree search method e.g., breadth-first search (BFS) or MCTS, to adapt these methods to utilize language models (LMs), we simply need to define how new states (successors) are generated and evaluated.
Action generation.
Given a trajectory of $i+1$ states during the search process, candidate actions $a^{(j)}_{i}$ in the action space $\mathcal{A}$ are sampled using an LLM-based policy $\pi_{\theta}$:
$a^{(j)}_{i}\sim\pi_{\theta}(a_{i}|x,s_{0},\ldots,s_{i}),\forall j\in\{1,\ldots,B\}$ (1)
where $B$ is the branching factor or the number of specified candidate actions. These sampled actions constitute the set $A_{s_{i}}$. We denote the transition function as $T:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$, so for an action $a_{i}$, $s_{i+1}=T(s_{i},a_{i})$.
State evaluation.
A value $v_{s_{i}|x}$ and a rationale for the value $r_{s_{i}|x}$ of a state $s_{i}$ is generated using a LLM-based value model $V_{\phi}:\mathcal{X}\times\mathcal{S}^{*}\rightarrow\mathbb{R}\times\mathcal{L}$, where $\mathcal{S}^{*}$ is the set of all finite sequences over the state space $S$ and $\mathcal{L}$ is the space of natural language sequences:
$(r_{s_{i}|x},v_{s_{i}|x})\sim V_{\phi}(x,s_{0},\ldots,s_{i})$ (2)
Note that because $V_{\phi}$ is an LLM, it is usually prompted in a chain-of-thought wei2022chain manner to first generate $r_{s_{i}|x}$ and then conditionally generate $v_{s_{i}|x}$ based on the rationale. For notational simplicity, will denote these two generated entities as $r_{s_{i}|x}\sim\text{Rat}(V_{\phi}(x,s_{0},\ldots,s_{i}))$ and $v_{s_{i}|x}\sim\text{Val}(V_{\phi}(x,s_{0},\ldots,s_{i}))$. While previous work zhou2023lats ; koh2024tree ; yao2024tree ; yu2024exact do generate rationales during state evaluation, usually they are generated purely to leverage the performance improvements of LLMs when asked to rationalize and are subsequently discarded. As described in §3, inspired by self-taught reasoning zelikman2022star , we explicitly use these rationales during self-improvement by fine-tuning on them along with the value to learn how to better estimate state values in a given domain.
3 Better State-Value Estimation with Self-Taught Lookahead
In this section, we present our proposed self-taught lookahead method. §3.1 and §3.2 describe the STL improvement procedure, while §3.3 explains how a STL-improved value model operates during inference. See Figure 2 and Algorithm 1 in Appendix C for an overview of the method.
3.1 Generating Rollouts
STL assumes a static policy model $\pi_{\theta}$ and only trains the value model through one or more iterations of self-improvement. We denote the value model initialized with a base LLM $V_{{\phi}_{0}}$ and the value model used in a subsequent iteration $k$ to assign values and generate rationales $V_{\phi_{k}}$.
An iteration $k$ of STL starts with a dataset $\mathcal{D}_{\texttt{rollout}_{k}}\subset\mathcal{X}$ of natural language tasks for the current iteration. For each $x_{i}\in\mathcal{D}_{\texttt{rollout}_{k}}$, we roll out the search tree using $\pi_{\theta}$ and $V_{\phi_{k}}$. Using tree search enables us to collect a diverse set of states so that the value model trained on these states’ values can better generalize to unseen states and tasks. We demonstrate this generalization in §4. When visiting state $s_{j}$ on the trajectory $\{s_{0},\ldots,s_{j}\}$ during tree search, we compute $s_{j}$’s lookahead value, $y_{s_{j}}$:
$y_{s_{j}}\leftarrow\gamma\max_{a\in A_{s_{j}}}\Big\{\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big\}$ (3)
where $\gamma$ is the discount factor. Since tasks are episodic, we set $\gamma=1$ bernstein2001reinforcement ; van2005survey ; leffler2007efficient . These lookahead values capture a better estimate of the true value of $s_{j}$ as they account for $_{j}’s s$uccessor states. In §7, we describe how generating and learning from these $y$’s is similar to fitted value iteration.
Action-outcome rationales.
However, alone, these lookahead values fail to reflect why a given state is valuable as they do not capture (1) which action yielded the best (highest value) successor state and (2) why the best successor state was assigned a high value by $V_{\phi}$.
To better capture the state transition dynamics, we also generate action-outcome rationales when visiting a state $s_{j}$. These rationales are of the form “{action} {outcome_state} {value_rationale}” where action is selected by the $\max$ operator in Equation 3, outcome_state is the state observed after taking the action in the environment and value_rationale is the rationale for the evaluation of this successor state generated by $V_{\phi_{k}}$. Fine-tuning on these rationales will enable a value model to predict the result of taking an action and incorporate this prediction (lookahead) into the current state’s value estimate. Formally, we can define these action-outcome rationales $o_{s_{j}}$:
[…]
where $\cdot||\cdot$ denotes concatenation, $s^{*}_{j+1}\leftarrow T(s_{j},a^{*}_{j})$, and
$a^{*}_{j}\in\operatorname*{arg\,max}_{a\in A_{s_{j}}}\Big\{\text{Val}(V_{\phi_{k}}(x_{i},s_{0},\ldots,s_{j},T(s_{j},a)))\Big\}\\$ (5)
The training data set for iteration $k$ is thus a set of tuples: $\mathcal{D}_{k}=(s_{k},o_{s_{k}},y_{s_{k}})$. Depending on the task, it might be necessary to automatically filter out tuples that have malformed rationales or account for the same state seen multiple times in different iterations (see Appendices DE, and F).
3.2 Fine-Tuning the Value Model
We start training the new value model $V_{\phi_{k+1}}$ from the initial or base LLM value model $V_{\phi_{0}}$. We can then train using standard fine-tuning negative log-likelihood loss for the generation of both the action-outcome rationale and the lookahead value ($o_{s}||y_{s}$) of the state. We train the value model to generate the rationale before estimating the value. Automatically constructed text or formatting, as seen in Figure 2, is applied for easier learning.
3.3 Search after Self-Taught Lookahead
A value model resulting from iteration $k$ of STL ($V_{\phi_{k+1}}$) can directly replace a value model in any search algorithm, such as Greedy Search (§4.1 and §4.2) and BFS (§4.3). As shown in Figure 2 , $V_{\phi_{k+1}}$ simulates a step of lookahead for the state $s_{n}$ i.e. for $({r}_{s_{n}|x},{v}_{s_{n}|x})\sim V_{\phi_{k+1}}(x,s_{0},\ldots,s_{n})$,
[…]
where $\tilde{a}_{{n+1}}||\tilde{s}_{{n+1}}$ is a simulated lookahead step and $\tilde{r}_{s_{n+1}|x}$ is its value rationale.
4 Experiments
We benchmark our proposed STL self-improvement approach on applied web agent tasks, multi-step question answering, and math puzzle tasks [^1] Our code is available at https://github.com/ethanm88/self-taught-lookahead..
4.1 Web Tasks
As mentioned in §1, it is particularly challenging and expensive to gather ground truth web task completion data zhang2024rest_mcts . To benchmark our STL method on web tasks, we utilize WebShop yao2022webshop , which consists of interactive web tasks involving searching for and purchasing an item that matches a short natural language specification. This benchmark is an ideal test bed to demonstrate the ability of our approach, as, unlike other web task datasets zhou2023webarena ; koh2024visualwebarena , ground truth reward is provided for all tasks, allowing a direct comparison between STL and methods in the Reward and Demo Learning and Reward-Guided Inference settings which use this reward.
Table 1: Score and success rate (SR) on WebShop. Results marked with $\dagger$}} are taken from previous work zhou2023lats ; putta2024agent . Value functions marked with $\ddagger$} are fine-tuned. We observe a near 40% improvement in success rate when using the STL value function compared to the llama-3.1-8b-instruct base value model in the greedy setting. We compute statistical significance of Reward and Demo Free methods against the underlined results ($^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001)$ using the paired bootstrap test berg2012empirical . Best results in the Reward and Demo Free setting are bolded.
Setting Method Policy Value Mini Test Set (50) Full Test Set (500)
Score $\uparrow$ SR $\uparrow$ Score $\uparrow$ SR $\uparrow$
[…]
Reward and Demo Learning
AgentQ
[…]
—— —— —— —— 50.5
Reflexion gpt-3.5-turbo —— 77.2 46.0 72.9 41.3
Reward-Guided Inference LATS gpt-3.5-turbo gpt-3.5-turbo 75.9 38.0 —— ——
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 70.0 26.0 67.7 26.4
gpt-3.5-turbo r1-distill-llama-8b 68.4 24.0 66.3 24.6
gpt-3.5-turbo gpt-3.5-turbo 71.5 38.0*** 70.6*** 35.6***
gpt-3.5-turbo gpt-4o 72.9* 42.0*** 71.5*** 40.6***
gpt-4o llama-3.1-8b-instruct 71.6 28.0 67.2 25.8
gpt-4o r1-distill-llama-8b 71.6 32.0* 66.5 25.6
gpt-4o gpt-3.5-turbo 77.4*** 46.0*** 72.4*** 38.8***
gpt-4o gpt-4o 74.4** 46.0*** 71.4*** 40.8***
MCTS Baseline gpt-3.5-turbo llama-3.1-8b-instruct 71.9 34.0** —— ——
[…]
Human Expert —— —— —— 76.1 54.0 82.1 59.6
STL for web tasks.
Following the empirical advantages on agent tasks identified by yu2024exact , we generate training data with MCTS by performing a step of lookahead at each step during rollout. Note that we use the LLM value model value outputs as a proxy reward to guide UCT (Upper Confidence bounds applied to Trees) kocsis2006bandit selection like yu2024exact , instead of ground truth reward zhou2023lats . We perform STL with a gpt-3.5-turbo gpt-35 policy to be consistent with previous work zhou2023lats ; shinn2024reflexion as well as a gpt-4o GPT-4o policy and fine-tune a llama-3.1-8b-instruct dubey2024llama value function. STL is performed by rolling out 50 tasks from the WebShop training set, resulting in 1161 training examples, which we find is sufficient for significant performance improvement. We find that training a separate value model at each depth allows us to train with smaller LLMs with fewer active parameters during rollouts (see Appendix D.3 for more details). Also, while we perform data generation during STL with MCTS, we evaluate the agent using the trained value models with greedy search, where the next action is greedily chosen based on the values of the policy’s proposed actions (see Algorithm 2 for more details). Finally, we find that a single iteration of STL is sufficient to see significant improvement over using a base LLM-initialized value model, and that multiple iterations do not yield additional performance improvements due to difficulties in simulating more than one step ahead, given the complexity of the environment. This conclusion has been corroborated by other LLM agent works gu2024your ; chae2024web . However, in §4.3 we show that this multi-step simulation is possible for simpler tasks. See Appendix D for further details about data generation, training, and simulation.
Baselines.
Within Reward and Demo Learning approaches, we include IL and RL methods originally proposed in the WebShop work that train BERT devlin2019bert and BART lewis2019bart models on human demonstrations and ground truth reward yao2022webshop . We additionally compare to the current state-of-the-art approach AgentQ putta2024agent , which finetunes a larger xLAM-v0.1-r-46.7b policy on rolled out MCTS search trees using direct preference optimization rafailov2023direct . We also include Reward-Guided Inference methods such as Reflexion shinn2024reflexion and LATS zhou2023lats , which work by prompting closed-source LLMs within a framework, e.g., MCTS, that is guided by ground truth reward. Finally, in the Reward and Demo Free setting, we include greedy search and MCTS [^2] We use LLM value as a proxy reward to guide UCT, like in the data generation phase of STL baselines with a ReACT yao2022react prompted base LLM (llama-3.1-8b-instruct, gpt-3.5-turbo, gpt-4o, r1-distill-llama-8b guo2025deepseek ). All value models are prompted with few-shot examples and are asked to provide reasoning before a numerical value. Following LATS for a fair comparison, we use a branching factor of 5 for all methods and 30 iterations for MCTS-based approaches. This makes LATS and MCTS baselines strictly more computationally expensive than greedy methods. We use the pass@$3$ chen2021passk for methods that do not have access to reward at inference time. Finally, we include a comparison to human expert performance measured in the original WebShop paper yao2022webshop .
Results and discussion.
The WebShop average reward (Score) and success rate (SR) of evaluated methods are presented in Table 1, with additional baselines in Table 4. We present results on both the full WebShop test set and on the mini test set of 50 tasks used by zhou2023lats , as we find running LATS and other MCTS methods on the entire test set is computationally expensive. Both of these sets are distinct from those seen during STL. Within the Reward and Demo Free setting, we find that STL matches the performance of using a gpt-4o value model and leads to a greater than 7% improvement in average reward and a 39% improvement in success rate relative to a base llama-3.1-8b-instruct value model, both of which are statistically significant improvements ($p<0.001$) using the paired bootstrap test berg2012empirical . Moreover, we find that STL even performs similarly to Reward-Guided Inference methods that have access to ground truth rewards. Lastly, we find no statistically significant difference between using a gpt-3.5-turbo or gpt-4o policy due to high action diversity in our setup (see Appendix D.2); thus, we evaluate search methods with a single policy (either gpt-3.5-turbo or gpt-4o) in subsequent experiments.
Table 2: Ablation study on the impact of fine-tuning with different combinations of information from lookahead, namely lookahead values (lv), textual representation of the next best action and successor state (tr), and the value rationale for the successor state (r). The underlined results are from the base model before any fine-tuning.
Fine-tuning Data Setup Score $\uparrow$
llama-3.1-8b-instruct 70.0
+ lv 76.0
+ lv + tr 74.4
+ lv + tr + r (stl) 78.3
Reasoning ablation.
In the Reward and Demo Free setting, we also perform ablations on the set of information from the step of lookahead used to fine-tune the value model during self-improvement. Specifically, we compare STL to variants that use only subsets of the information derived from lookahead. As mentioned in §3, this information includes (1) the lookahead value, (2) the textual representation of the next-best action and its successor state, and (3) the value rationale for the successor state. The results of this ablation in Table 2 demonstrate that regressing solely on lookahead values and further incorporating state transitions from lookahead does improve performance relative to the base model. However, learning also from the value rationale of the successor state, as done in STL, yields additional performance gains over these other settings. These results substantiate the claims made in §3 about the necessity of learning from action-outcome rationales, a key difference between STL and both classical RL gordon1999fitted_value_iteration and other LLM tree search works feng2023alphazero ; zhang2024rest_mcts which fine-tune an LLM value model on numerical values only. Note for fairness, we modify the loss-masking to improve the lookahead value-only baseline and include this higher result in the table. See Appendix F.4 for more details about loss masking and this baseline.
4.2 Multi-Hop Question Answering
We also investigate the efficacy of STL on applied reasoning for retrieval-based question-answering. We specifically utilize the HotpotQA yang-etal-2018-hotpotqa benchmark, consisting of multi-hop question-answering tasks that require retrieving and reasoning over multiple Wikipedia entries. We use the same setup as in §4.1 to generate data by rolling out with MCTS, but instead roll out 500 tasks from the training dataset since actions (search terms) proposed by the policy lack diversity compared to web tasks.
Baselines.
In the Reward and Demo Learning setting, we evaluate on R1-Searcher song2025r1 , which performs RL on outcome-based answer correctness rewards on the multi-step retrieval QA task. As in §4.1, in the Reward-Guided Inference setting, we include Reflexion and LATS baselines, which have access to ground truth correctness during search. Among Reward and Demo Free methods, we compare STL to greedy search with closed-source models.
Results and discussion.
Table 3 presents the answer match rate of all evaluated methods. Following feng2024agile , we prompt gpt-4o to help evaluate answer correctness. The prompt used is in Appendix E.3. This fuzzy matching correctness is exposed as the reward for Reflexion and LATS, but not to the greedy methods in the Reward and Demo Free setting. We evaluate on a set of 500 unseen questions and also a smaller set of 50 examples due to the high cost of LATS. STL on llama-3.1-8b-instruct outperforms a gpt-4o value model and the R1-Searcher method while approaching the performance of Reward-Guided Inference methods, which can verify predicted answer correctness during inference. Of the Reward and Demo Free approaches, STL is the only method that is statistically significant compared to the llama-3.1-8b-instruct baseline ($p<0.001$).
Table 3: Match rates on HotpotQA. Value functions marked with $\ddagger$}} are fine-tuned. Statistical significance with the paired bootstrap test of Reward and Demo Free methods against the underlined results ($^{***}p<0.001$) is provided. The best  Reward and Demo Free results are bolded.
Setting Method Policy Value Match Rate $\uparrow$
Test Set (50) Test Set (500)
[…]
Reflexion gpt-3.5-turbo 70.0 66
Reward-Guided Inference LATS gpt-3.5-turbo gpt-3.5-turbo 70.0
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 60.0 56.4
gpt-3.5-turbo gpt-3.5-turbo 62.0 56.0
gpt-3.5-turbo gpt-4o 68.0 57.6
Reward and Demo Free
Greedy w/
STL (Ours)
[…]
4.3 Math Puzzles
Figure 3: BFS Game-of-24 performance on tasks seen and unseen during STL.
Finally, we also study the performance of STL on the Game-of-24 task yao2024tree , where the goal is to construct a mathematical expression with 4 provided integers to obtain 24. The task was originally for Reward and Demo Free methods, so it serves as a good benchmark for STL.
STL for Game-of-24.
For this task, we generate data using breadth-first search (BFS) rather than MCTS to be consistent with the original Tree-of-Thoughts yao2024tree approach. We use a gpt-4o policy and a llama-3.1-8b-instruct base value function in the first iteration. As described in §3, we replace this base value function with a trained model in each subsequent iteration. STL is run for four iterations of 25 puzzles. For this task, we do not have explicit environment observations, but instead use the policy’s arithmetic to combine two numbers as a pseudo-observation. See Appendix F for further details.
Baselines.
We compare the performance of value models learned via STL with Reward and Demo Free BFS baselines that use the same gpt-4o policy. Specifically, we experiment with a gpt-4o value model and an algorithmic oracle evaluator chen2024search_useful . This oracle runs a recursive algorithm to verify whether the current state (set of numbers) can be combined to reach $24$. Search performance with this oracle is an upper bound on the performance improvement possible from improving the value function under a static policy.
Results and discussion.
Figure 3 shows the performance of evaluated methods on a set of 50 tasks seen during STL and a set of 50 more challenging (determined by lower human solve percentages), unseen tasks. On both sets, STL matches or outperforms a gpt-4o value model. However, STL’s performance on seen tasks monotonically increases for the first three iterations, while its performance on unseen tasks decreases before increasing. This phenomenon is due to the limited number of tasks the value model sees during training during the first couple of iterations. Specifically, if value models are not exposed to enough actions and their lookahead values during training, they fail to generalize well to unseen tasks. Limiting the number of tasks per iteration also limits the quantity and diversity of actions and values seen. A full analysis of this result can be found in Appendix F.5.
Figure 4: Compute and environmental efficiency during evaluation on WebShop with a gpt-3.5-turbo policy (left). Compute efficiency is measured in total (prompt and completion) tokens. Environmental efficiency is measured by the number of states expanded (webpages visited). The distribution of tokens (closed vs. open source models) used during search is also shown (right). Value models are specified in parentheses.
5 Efficiency Analysis
In this section, we compare the efficiency of STL with prior methods on WebShop. We study efficiency tradeoffs from two perspectives (1) model costs and (2) environment usage. Additionally, in §5.3 we explore how performance changes when scaling the size of the value model trained during STL.
5.1 Compute and Cost Efficiency
Keeping compute requirements and costs low is critical, especially for agents automating routine, repetitive tasks. Since STL can be used to improve an open-source LLM like llama-3.1-8b-instruct, we can transfer computation from more expensive closed-source models like gpt-4o to open-source alternatives while maintaining performance. Figure 4 (right) demonstrates this transfer, as STL uses more than 50% fewer tokens generated from closed-source than greedy search with a gpt-3.5-turbo or gpt-4o value model. We also compute the monetary costs of different methods. Unlike other methods, STL incurs costs for data generation and fine-tuning in addition to inference, but these are one-time costs that do not scale with agent use and are quite modest at \$8.54 and \$1.76, respectively. We plot inference costs against the average WebShop reward in Figure 5 (left), and find that STL is Pareto optimal, $23\times$ cheaper than MCTS methods like LATS, and $5\times$ cheaper than performing greedy search with a gpt-4o value model. See Appendix G for details about the cost calculation.
Figure 5: Tradeoff between performance and efficiency on WebShop with a gpt-3.5-turbo policy. Pareto frontiers of existing methods and baselines are shown, illustrating the optimality of STL when considering the tradeoff between inference cost and average reward (left) and between environmental usage and average reward (right). Reward-Guided Inference methods are presented in gold and not included in the Pareto frontier since they belong to a different information setting.
5.2 Environmental Usage
It is often crucial for an agent taking actions in physical or digital environments to be environmentally efficient or conservative in the number of states it visits while performing a task. In the case of digital web agents, taking many steps per task through exhaustive tree search may put an unnecessary burden on web servers, especially as agents are deployed at scale. Additionally, allowing web agents to search widely when equipped with personal information or the ability to make purchases may lead to unintended privacy disclosures or financial loss, respectively. In some environments, taking many actions may also lead to unreasonable task completion times. Figure 4 (left) and Figure 5 (right) present the environmental efficiency measured by the count of expanded states or visited sites in the WebShop environment. Considering WebShop score, STL is Pareto optimal and requires expanding half as many states as MCTS-based methods like LATS. Moreover, unlike LATS, STL does not require irreversible actions (actually clicking Buy Now on a product page) required to obtain reward.
5.3 STL Scaling Trends
Figure 6: STL scaling trends on WebShop for llama-3 and qwen-2.5 families with a gpt-3.5-turbo policy.
As 8B STL models can match the performance of a gpt-4o value model, is it possible to use even smaller models for STL while maintaining good performance? STL requires models to (1) provide generally consistent values out-of-the-box so that it is possible to compare successor states during data generation and (2) learn to generalize to unseen tasks and states, both of which may be challenging for smaller models. We explore this STL scaling trend on WebShop with $\leq$ 8 billion parameter models in the llama 3 family [^3] As the llama-3.1 family lacks smaller models, we use 1B and 3B models from the llama-3.2 family. dubey2024llama and $\leq$ 7 billion parameter models in the qwen-2.5-instruct family yang2024qwen2 . The results presented in Figure 6 demonstrate that while performance does generally decrease with fewer parameters, the smaller 3B parameter models in both families do approach gpt-4o performance. This result suggests that using smaller models for STL is feasible, making large-scale agent deployment to new domains more practical.
6 Discussion: When to use STL?
As discussed in §4 and further detailed in Appendix D.2, achieving strong performance with STL requires high action diversity (a large number of possible actions at each step). When action diversity is low, the effect of the value model on search performance is diminished, and it is necessary to roll out more tasks to obtain enough data to fine-tune the value model. For example, Section 4.2 notes that due to its low action diversity, HotpotQA shows smaller relative gains from STL compared to WebShop and also requires rollouts of 500 tasks (10 times more than WebShop) to get a large enough dataset for fine-tuning.
Additionally, STL requires later states to provide good value estimates that can be backed up and learned during training on lookahead results. Search performance will benefit most from STL on tasks where state transitions are consistent throughout the environment, i.e., the same or semantically similar actions yield similar outcomes. For instance, clicking the “Low to High” button on any search results page consistently orders items by price. This consistency enables the STL value model to accurately simulate a step of lookahead (§3.3), leading to better state-value estimation and improved downstream search performance. Tasks that have stochastic transitions or have inconsistent transitions when actions are semantically similar may not show large improvements with STL. Fortunately, most popular tasks where LLM agents have been deployed, such as web navigation, have deterministic and fairly consistent transitions.
7 Related Work
Classical RL.
STL is loosely inspired by fitted value iteration (FVI) gordon1999fitted_value_iteration , which generalized value iteration bellman1957dynamicprog to the tabular setting. In an iteration of FVI, target values are computed using the Bellman update and used to train a new value model from the previous model checkpoint using least squares regression. The iterated values in FVI are computed similarly to the lookahead values $y_{s_{k}}$ in §3.1, but with STL, no ground truth reward is assumed, the value model is non-Markovian, and actions are deterministic. Instead of learning directly from iterated values, with STL, they are concatenated with action-outcome rationales, and together, these sequences are used to fine-tune the LLM value model from scratch at each iteration rather than from the previous model checkpoint as in FVI.
LLM self-improvement.
A variety of previous work has shown that LLMs can self-improve with iterative prompting techniques huang2022large ; weng2022large ; madaan2024self and have applied these methods to various domains, from agents kim2024language to privacy protection chen2023can . A separate line of work focuses on bootstrapping a small training dataset through a self-training process to improve either the reasoning policy model gulcehre2023rest ; singh2023rest_em ; jung2023impossible or the verification or reward model hosseini2024v using synthetically generated data. While most self-training approaches utilize outcome-based reward models, other work aksitov2023rest_react ; zhang2024rest_mcts derive process-based rewards like STL to evaluate each step in the reasoning chain.
Training reasoning agents.
The majority of prior work on training reasoning agents focuses on performing SFT on human-annotated trajectories yao2022webshop ; lai2024autowebglm , synthetically generated trajectories chen2023fireact ; furuta2023multimodal ; zhang2024xlam ; liu2024apigen ; patel2024large ; murty2024nnetscape ; ou2024synatra ; su2025learn , or a combination of both zeng2023agenttuning ; zhang2024agentohana . Other work has trained agents from tree search-generated data gandhi2024stream ; putta2024agent ; zhang2024rest_mcts or with explicit reinforcement learning  peiyuan2024agile ; wei2025webagent ; feng2025group , but these methods usually require ground truth reward. While prior work has explored self-improving reasoning agents, these approaches fail to generalize beyond the instructions encountered during self-improvement patel2024large or require fine-tuning frontier models like gpt-4o to achieve generalization yu2024exact .
8 Conclusion
We propose STL as an efficient method to improve the value model employed during search. This efficiency primarily stems from STL’s design for information-scarce settings, where models learn from state-transition dynamics. Additionally, because STL enables self-improvement on small models deployable with less exhaustive search, it yields significant reductions in both computational cost and environmental usage. Therefore, the STL framework could help enable the more realistic learning and deployment of agent systems.
Acknowledgments
We would like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This research is supported in part by the NSF under grant number IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. We also appreciate Jungsoo Park, Yao Dou, Geyang Guo, and the anonymous NeurIPS reviewers for their valuable feedback, which helped to improve the paper.
Appendix A Broader Impacts
Our STL method enables LLMs to self-improve at search by capturing the mechanics of traditional RL algorithms in natural language. While we leverage this technique for better state-value estimation, it can likely be applied to other problems. Additionally, we show in §5, that STL requires significantly less compute overhead than similarly performing methods, thus reducing the energy consumption required for search during inference, helping to improve the sustainability of agent deployment.
As with all self-improvement work [45, 66], having models self-improve without human supervision may enable agents to learn how to take actions towards task completion that are not well-aligned with human values or preferences. To prevent misuse, we only release code to reproduce experiments in this paper and not to improve general-purpose agents for which these risks might be more prevalent. While these harms are out of the scope of this work, we encourage future research in this area.
Finally, we note that it is possible to use our value model as a sort of policy to directly select actions in environments with a finite set of actions from each state. While we did not explore how well such a policy would work, improving value estimation to improve the policy may be an interesting direction for future work.
Appendix B Limitations
During self-improvement, STL does require task specifications like “I am looking for a queen-sized bed that is black …” (see Figure 2) for web tasks. However, this assumption is well-founded as prior self-improvement work, such as that on LLM alignment [67, 59], also assumes tasks, or in their case, user prompts, are provided to initialize the self-improvement process. Additionally, as mentioned throughout the paper, this setting is much more reasonable than Reward and Demo Learning and Reward-Guided Inference settings, which require ground truth reward and/or human demonstrations along with task specifications.
Additionally, due to compute constraints, our experiments utilize models up to 8 billion parameters for STL. However, our results in §5 demonstrate a scaling trend in performance with respect to model size, indicating that using larger models may yield more performant agents. Given these positive results, we leave applying STL to larger models to teams with larger resource budgets.
STL requires actions to be grounded in language to be able to leverage LLMs’ strong priors. Therefore, STL may struggle in traditional RL tasks like gridworlds, which may have a large state and action space and transitions that are not grounded in language. However, prior work has contextualized gridworld environments by adding landmarks [11], so similar techniques may be used to obtain the necessary consistent and grounded transitions for STL to be successful.
Algorithm 1 Self-Taught Lookahead
1:Set of tasks $\mathcal{D}_{\texttt{rollout}}$, Base LLM $M$, num iterations $n$, num tasks per iteration $m$
[…]
5:for $k\leftarrow 1$ to $n$ do
6:  $\mathcal{D}_{\texttt{rollout}_{k}}\leftarrow\mathcal{D}_{\texttt{rollout}}[m\cdot(k-1):m\cdot k]$ $\triangleright$ Select tasks for iteration $k$
7:  for $x\in\mathcal{D}_{\texttt{rollout}_{k}}$ do
8:   $T\leftarrow\texttt{rollout\_tree}(\pi_{\theta},V_{\phi_{k-1}},x)$ $\triangleright$ Generate rollout search tree for task $x$
9:   $y\leftarrow\texttt{calculate\_lookahead\_values}(T)$
10:   $o\leftarrow\texttt{calculate\_action\_outcomes\_rationales}(T)$
11:   $y_{\texttt{filtered}},o_{\texttt{filtered}}\leftarrow\texttt{task\_specific\_filter}(y,o)$ $\triangleright$ Apply task-specific filtering if applicable
12:   $\mathcal{D}_{\texttt{Train}_{k}}\leftarrow\texttt{add\_new\_data}(\mathcal{D}_{\texttt{Train}_{k}},(y_{\texttt{filtered}},o_{\texttt{filtered}}))$
13:  end for
14:  $V_{\phi_{k}}\leftarrow\texttt{fine\_tune}(V_{\phi_{0}},\mathcal{D}_{\texttt{Train}_{k}})$ $\triangleright$ Finetune from base model
15:end for
Appendix C Algorithms
The STL algorithm is presented in full in Algorithm 1. For information about the task_specific_filter, see Appendix D.3 and Appendix F.4.
Additionally, the algorithm for greedy search as used in §4 is presented in Algorithm 2.
Appendix D Self-Taught Lookahead on WebShop
In this section, we outline information about WebShop [63] and the implementation details of running STL on the benchmark.
D.1 WebShop Task
There are two main types of actions in the WebShop task:
  • search[query]:
    Search actions allow the user to search for a particular item with a natural language query, e.g., search[easy to use medium color face kit less than 40 dollars]. This action can only be taken on the search page, which is also the initial / home page of the WebShop interface.
  • click[button]: Click actions are discrete actions but can take many forms, which we enumerate below:
    • click[product]: to select a relevant product from the search results e.g. click[B09B6SH764] where B09B6SH764 is a product code.
    • click[attribute]: to toggle on an attribute or option on the product page of an item, e.g. click[small]
    • click[Buy Now]: to buy the selected item - this is a terminal action that yields the ground truth reward. This action is not allowed to be taken in STL search but is allowed in other search, RL, and prompting methods [63, 49, 74].
    • Other navigation buttons: other navigation buttons include click[Back to Search], click[<Prev], click[Next>], click[Description], click[Features]. To simplify trajectories, we generally restrict the ability for models to take these actions in all settings following [63].
WebShop provides a textual representation of webpages in simple mode. An example of this representation for search results is shown in Figure 7.
Algorithm 2 Greedy Search
1:LLM policy $\pi_{\theta}(\cdot)$, LLM value model $V_{\phi}(\cdot)$, Initial state $s_{o}$
[…]
3:while $s_{i}$ is not terminal do
4:  $a_{i}\leftarrow\operatorname*{arg\,max}_{a\in A_{s_{i}}}\Big\{\text{Val}(V_{\phi}(x,s_{0},\ldots,T(s_{i},a))\Big\}$ $\triangleright$ Greedily pick best action
5:  $s_{i+1}\leftarrow T(s_{i},a_{i})$
6:  $i\leftarrow i+1$
7:end while
WebShop Textual Representation [Back to Search] Page 1 (Total results: 50) [Next >] [B0972Q1T8T] Cosycost USB Microphone, Condenser Computer PC Gaming Microphone for PS4/5 Laptop Windows Mac OS Android Phone, Noise Cancelling Instant Mute, Studio Mic for Voice, Music Recording, Podcasting, Streaming \$32.99 [B09N3M6H2Z] Wired Stereo Headset Noise Cancelling Microphone with in-line Controls/Volume Controller, All-Day Comfort Design, Works for Playstation, Nintendo Switch, PC with USB Connection (HS-HP101UNCBK) \$199.99 [B072L2D6LY] Andrea Communications NC-255VM USB On-Ear Stereo USB Computer Headset with Noise-Canceling Microphone, in-Line Volume/Mute Controls, and Plug \$34.59 [B071H84LTJ] Andrea Communications NC-455VM USB Over-Ear Circumaural Stereo USB Computer Headset with Noise-Canceling Microphone, in-Line Volume/Mute Controls, and Plug \$49.24 [B08GLJSWJ9] Jiade USB Headset with Noise Canceling Microphone for CallCenter Skype Chat, Computer Phone Headset Voice Recognition Speech Dictation, PC Headphone with Mic Mute Volume Control Binaural Golden \$9.99
Figure 7: Example of simple mode textual representation of the state with the WebShop benchmark.
D.2 Prompts
The prompt used to generate actions in WebShop is presented in Figure 8. Notice that we do not use think actions part of the classical ReACT framework [65] like [63] or [74] because evaluating the value of these actions is difficult as they have no observation. Instead, we prompt the policy model to provide a rationale while generating possible actions. Also, note that we prompt the policy multiple times, adding to the list of actions that are not allowed and removing from the list of actions that are allowed. This change to a “selection” policy enables action diversity, which we find is otherwise low even with prompting the policy at high temperature. This change also likely explains why there is no statistically significant difference between using a gpt-3.5-turbo and gpt-4o policy in §4.1.
Likewise, the prompt used to evaluate states is presented in Figure 9. Note that this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. We note that the Likert scale used was crucial to obtaining consistent value outputs on which we could perform STL. We also use a special value estimation prompt whenever on a product listing page to select attributes (see Figure 10) in order to obtain consistent values. Specifically, we convert the 4-point scale into a -2 to 2 scale and add the value to the value of the prior state’s (product selection action) value. We tried to use this in our baselines, but we found that it actually harmed the baseline scores. However, in exploratory experiments, it helped yield more consistent values, so we use this for STL.
For all value estimates (base model or fine-tuned), we prompt the value model 5 times and use the average score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 5 scores. After the rebuttal period, we found an instruction to further discount values in the prompts to the fine-tuned models. While we could not rerun all results due to inference costs, we reran the STL evaluation on the mini test set and found results were maintained.
In total, the data generation phase on WebShop yielded a total of 1162 examples, collected from rolling out search trees for 50 tasks. An example of the rationale structure of the training data is presented in Figure 11.
WebShop Generation Prompt You are a web agent, select the best next action for the search to fulfill the task. Example tasks are shown below. Provide a rationale for your selection BEFORE you provide the action. NOTE: You can only select actions that are provided in the Possible Actions list. You MAY NOT select actions in the Not Allowed list. NOTE: You must output BOTH a rationale and an action. NOTE: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page. Example Tasks: {few shot examples} —————————————————————————————————– New Task: {task} Actions Not Allowed: {not_allowed_actions} Possible Next Actions (REMINDER: You can only select actions from this list.): {possible_actions} REMINDER: Do not select any of the following actions: ’Back to Search’, ’Next >’, ’< Prev’, ’Attributes’, ’Description’, ’Features’, ’Reviews’, even if they are available on the page.
Figure 8: Generation prompt for WebShop policy.
WebShop Value Estimation Prompt Given an item to purchase and a trajectory that aims to buy an item that exactly matches the specification, analyze how well the last action and observation align with the task. Provide a reflection that concludes with. "Thus the correctness score is s", where s is either 1, 2, 4, 6, 8, or 10. Use the following scale for scoring: 1: The last action and observed state is entirely irrelevant to the task or captures a purchase of an item that is completely unrelated to the specifications. 2: The last action and observed state captures a step with a low likelihood of leading to purchasing the correct item. 4: The last action and observed state captures a step with a moderate likelihood of leading to purchasing the correct item. 6: The last action and observed state captures a step with a high likelihood of leading to purchasing the correct item. 8: The last action and observed state captures a step with a very high likelihood of leading to purchasing the correct item. 10: The last action and observed state captures a step that will certainly lead to purchasing the correct item. Keep reflections short (<100 words). Follow the format of the rationale from the below example task. NOTE: the observation from clicking on the item will be the item’s product detail page. For instance, click[B078GWRC1J] will show the product detail page for the item with code B078GWRC1J which will include the item’s name (e.g. Bright Citrus Deodorant by Earth Mama), price (\$10.99), and other relevant details as well as options. NOTE: Assume none of the attributes on the product page are selected only provide the reflection for the last action. Example Tasks: {few shot examples} —————————————————————————————————– New Task: Respond with the reflection for the last observation of the new task ONLY. As a reminder the last action and observation is as follows: {last_action} Your response should start with "Reflection:" and end with "Thus the correctness score is …".
Figure 9: Value Estimation prompt for WebShop. This prompt was only used to prompt base models.
WebShop Attribute Prompt Given an item to purchase and a trajectory that aims to buy an item that exactly matches the specification, analyze how well the last action and observation align with the task. All the last actions you see will be selecting an attribute on the product page of a candidate item. Provide a reflection that concludes with "Thus the correctness score is s", where s is either 1, 2, 3, or 4. Use the following scale for scoring: 1: The attribute selected is opposite to the specified attribute in the task. 2: The attribute selected is irrelevant to the specified attribute in the task. 3: The attribute selected is an attribute mentioned in the instruction, but not all attributes mentioned are currently selected. 4: The attribute selected is an attribute mentioned in the instruction, and all attributes mentioned are currently selected. Keep reflections short (< 100 words). Follow the format of the rationale from the example task below. NOTE: When selecting attributes like click[x-large], click[red], etc., the observations will be “You have clicked x-large” or “You have clicked red” respectively. IMPORTANT: Assume options for size, weight, color, etc. on product pages are not selected unless explicitly mentioned in the trajectory (e.g., click[large], click[red], etc.). Selecting these attributes may be crucial to match the task requirements. Only provide the reflection for the last action. Example Tasks: {few shot examples}  New Task: {input} Respond with the reflection for the last observation of the new task ONLY. As a reminder, the last action and observation is as follows: {last_action} Your response should start with “Reflection:” and end with “Thus the correctness score is …”
Figure 10: Value estimation attribute prompt for WebShop. This prompt was only used to prompt base models during rollouts.
WebShop Rationale Structure Task and Current Trajectory:
Instruction:
i’m looking for a pair of women’s workout shorts with a drawstring waist. i need them to be extra large and in light gray, and price lower than 60.00 dollars [Search]
Action: search[women’s workout shorts drawstring waist extra large light gray < 60 dollars] Observation: removed for brevity Current Action: click[B09T3PJM1R] Current Observation: removed for brevity
STL Output:
I will evaluate the best successor state from the current state:
Best Next Action: click[x02c-gray] Observation of Best Successor State: You have clicked x02c-gray. Reflection of the Best Successor State: The last action selects the color ’x02c-gray’ for the item B09T3PJM1R. Based on the observation, this product’s color is indeed gray, which matches the specified criteria. Therefore, this product matches one of the attributes mentioned in the task. The last action and observation thus capture a step that selects an attribute mentioned in the instruction, but not all attributes mentioned (specifically the size attribute) are currently selected. Thus, the correctness score is 6.00 / 10.00.
Figure 11: Example of rationale structure from the STL training data rolled out on the WebShop task.
D.3 Implementing STL
With web tasks, the position of an action in a trajectory may influence its value. For instance, selecting a certain item I from search results early in the trajectory should have a higher value than selecting I a second time in the same trajectory. To account for this difference, we train value models at each position (depth) in the trajectory. Specifically, we limit trajectories to five steps and train four value models depths 1 to 4, only allowing a terminating Buy action on the final step.
We also filter out malformed rationales from the training data. Specifically, we remove rationales that do not provide the proper format, e.g., it does not exactly contain scaffolding like “Thus the correctness score is”.
Additionally, we generate lookahead rollouts using 5 iterations of MCTS, which, in practice, we find is sufficient for collecting diverse training data.
Finally, we compute loss over both input and output tokens to enable the value model to more quickly capture the dynamics of the environment from the trajectory. However, as we mentioned in §4.1, this practice harms the performance in the setting where the model regresses only on the lookahead value. Therefore, we found that when we only compute loss on the output tokens as in the normal SFT setting, the WebShop score in this setting increases from $70.9$ to $76.0$. This score is actually higher than the setting where we predict the future next best state without a rationale, which indicates that rationales may be key when we try to learn from transition dynamics. We are not sure why this is the case, and why there is not a larger gap between STL and this lookahead value baseline, but we believe that it could be due to small differences in the prompt due to different information settings and/or the prediction or future state potentially harming value estimation for clearly good or bad states.
D.4 Difficulty in Performing STL for Multiple Iterations.
Empirically, we find that STL after a second iteration on WebShop has a lower performance (average reward of 68.6, and success rate of 26.0) than after a single iteration. From a manual inspection of the lookahead and rationales generated, we notice that the second step simulated by the value model often does not match the true environment.
Table 4: Additional baselines on WebShop. See Table 1 for the full results.
Setting Method Policy Value Mini Test Set (50)
Score $\uparrow$ SR $\uparrow$
ReACT (No Search) gpt-3.5-turbo 68.9 36.0
gpt-3.5-turbo 70.0 36.0
Reward and Demo Free Greedy Baseline gpt-3.5-turbo qwen3-8b 68.6 34.0
Table 5: Score and success rate (SR) on WebShop with error bars. The full results are in Table 1.
Method Policy Value Full Test Set (500)
Score $\uparrow$ SR $\uparrow$
Greedy Baseline gpt-3.5-turbo llama-3.1-8b-instruct 67.7 ± 2.26 26.4 ± 3.86
gpt-3.5-turbo r1-distill-llama-8b 66.3 ± 2.30 24.6 ± 3.78
gpt-3.5-turbo gpt-3.5-turbo 70.6 ± 2.43 35.6 ± 4.23
gpt-3.5-turbo gpt-4o 71.5 ± 2.51 40.6 ± 4.33
gpt-4o llama-3.1-8b-instruct 67.2 ± 2.30 25.8 ± 3.84
gpt-4o r1-distill-llama-8b 66.5 ± 2.33 25.6 ± 3.83
gpt-4o gpt-3.5-turbo 72.4 ± 2.40 38.8 ± 4.29
gpt-4o gpt-4o 71.4 ± 2.49 40.8 ± 4.35
[…]
Human Expert —— —— 82.1 59.6
D.5 Additional Results on WebShop
We have included performance on other baselines in Table 4, including ReACT [65] baselines with OpenAI models and an additional greedy search baseline with a qwen3-8b reasoning value model. Results with error bars are also provided in Table 5.
D.6 Reflexion Performance on WebShop
Tracing bugs in Reflexion implementation
We originally ran Reflexion baselines on WebShop using the official Reflexion GitHub repository [^4] https://github.com/noahshinn/reflexion with three iterations, changing only the model from text-davinci-003 (a deprecated text completion model used in the original paper) to gpt-3.5-turbo. During the author rebuttal period, we found the following:
  • Other researchers had obtained similar Webshop performance to our initial Reflexion results ( 15%) using the unchanged official implementation with gpt-3.5-turbo (see GitHub issue #49).
  • Others also noted that changes to the prompts were needed to adapt the framework to conversational models to see improved performance in the 30-40% range (see GitHub issue #48).
  • Another researcher also identified a bug in the WebShop implementation that prevented the use of memory from prior iterations (see GitHub issue #36). In the discussion of this issue, the first author of the Reflexion paper acknowledged that this bug may have caused the lack of improvement of the Reflexion agent on WebShop that was reported in Appendix B.1 of the original Reflexion paper [49].
After modifying the prompts and patching this memory bug, the ReAct success rate on the 50-task test set is 36%, and the Reflexion success rate is 46% after three iterations, which is similar to the performance reported in other work [73]. We hope that this debugging process can be of some use to the community.
Reflexion vs. LATS.
In Table 1, Reflexion outperforms LATS, which is potentially unexpected as LATS is more computationally expensive than Reflexion. However, due to differences in the mechanisms of these two methods, it is not necessarily the case that LATS performance is lower-bounded by Reflexion performance. For instance, Reflexion enables the LLM policy to make wholesale changes to any steps in its trajectory after reflecting on previous failures. On the other hand, while LATS provides its policy and value LLMs with reflections on previous trajectories, it still relies on the traditional (non-neural) mechanics of MCTS. For example, LLM reflection is not involved during node selection, and instead, the traditional Upper Confidence bounds applied to Trees (UCT) heuristic is used.
Appendix E Self-Taught Lookahead on HotpotQA
E.1 HotpotQA Task
The HotpotQA [62] is a multi-hop question answering benchmark where correct answers require reasoning over multiple Wikipedia entries. There are three possible actions at each step:
  • search[entry]:
    Search provides the first five sentences of the corresponding Wikipedia entry if it exists, or provides five alternative existing entries.
  • lookup[string]:
    Lookup returns the next sentence in the entry containing the specified string.
  • finish[string]:
    Finish signifies the completion of the reasoning process, where the answer is specified with the provided string.
E.2 Prompts
The prompt used to generate actions for the HotpotQA task is presented in Figure 14. Likewise, the prompt used to evaluate states is presented in Figure 13. Many of the prompting details, such as the use of a disallowed action list, are similar to the WebShop task. However, one difference is that if five unique actions are not found due to a search term not being an entry, we add the top-5 similar terms returned by the retrieval model as possible actions.
In total, the data generation phase on HotpotQA yielded a total of 2708 examples, collected from rolling out search trees for 500 tasks.
E.3 Evaluation
We use the same gpt-4o evaluation prompt as [15] which is provided in Figure 13. Note that a match is computed when either there is an exact match or a match indicated by the gpt-4o evaluator in order to prevent any mistakes by the evaluator, such as indicating a non-match even though the two answers were an exact string match.
Additionally, unlike WebShop, we limit trajectories to a depth of four, and explicitly prompt models to provide a final answer on the fourth step in the trajectory.
E.4 Implementing STL
The implementation details of STL for HotpotQA are similar to WebShop in the training of multiple models at different depths, the filtering out of malformed rationales, and the use of 5 iterations of MCTS during data generation. However, one difference is that we feed the possible next actions to the value model so that it can provide coherent lookahead simulations and evaluations.
HotpotQA Generation Prompt Solve a question answering task with interleaving Thought, Action, Observation steps. Thought can reason about the current situation, and Action can be three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. After each observation, provide the next Thought and next Action. NOTE: You MAY NOT select actions in the Actions Not Allowed list. You have to change the wording of the query or lookup somehow. NOTE: Keep search queries and lookups short and concise since longer queries will not return any result. For instance, instead of searching Search[eastern sector of the Colorado orogeny], search Search[Colorado orogeny] and then Lookup[eastern sector]. Here are some examples: {few shot examples} —————————————————————————————————- New Task: {task} Actions Not Allowed: {not_allowed_actions} Thought:
Figure 12: Generation prompt for the HotpotQA policy.
HotpotQA Value Estimation Prompt Given a question and a trajectory to answer the question, analyze how well the LAST ACTION in the trajectory contributes to finding the answer. Consider ONLY the last action. The trajectories are labeled by pairs of thoughts that can reason about the current situation and actions that can be of three types: (1) Search[entity], which searches the exact entity on Wikipedia and returns the first paragraph if it exists. If not, it will return some similar entities to search. (2) Lookup[keyword], which returns the next sentence containing keyword in the current passage. (3) Finish[answer], which returns the answer and finishes the task. Provide a reflection that concludes with “Thus the correctness score is s”, where s is either 1, 3, 5, 7, or 10. Use the following scale for scoring: 1: The action is completely irrelevant to answering the question or there is no relevant search result (“Could not find” is in the observation). 3: The action’s observation provides information only at the background level to answering the question. 5: The action’s observation provides information that makes a small step towards answering the question. 7: The action’s observation provides information that makes a key step towards answering the question. 10: The action’s observation provides the final piece of information needed to answer the question. Reminder: If you see “Could not find” in the observation, the correctness score is 1. Keep reflections short (< 100 words). Follow the following examples: {few shot examples} —————————————————————————————————- {input}
Figure 13: Value estimation prompt for HotpotQA. This prompt was only used to prompt base models.
HotpotQA Evaluation Prompt Based on the provided question and reference answer, please determine if the response is correct or incorrect. Begin by articulating your rationale, and conclude with a single word judgment: ‘Yes’ for correct or ‘No’ for incorrect. question: {question} reference answer: {reference} response: {response}
Figure 14: Evalaution prompt for the HotpotQA predicted answers.
Game-of-24 Generation Prompt Use numbers and basic arithmetic operations (+ - * /) to obtain 24. In each step, you are only allowed to choose two of the remaining numbers to obtain a new number. Follow the example format exactly. {few shot examples} —————————————————————————————————- {input}
Figure 15: Generation prompt for the Game-of-24 policy.
Game-of-24 Value Estimation Prompt Evaluate if the given numbers can reach 24 (sure/likely/impossible) Follow the example format exactly. Only evaluate the last example. {few shot examples} —————————————————————————————————- {input}
Figure 16: Value estimation prompt for Game-of-24. This prompt was only used to prompt base models.
Appendix F Self-Taught Lookahead on Game-of-24
F.1 Game-of-24 Task
As introduced by [64], the Game-of-24 is a mathematical reasoning task that involves combining four numbers e.g. “2 3 4 5” together with mathematical operations i.e. ++, -, //, $\times$ in order to obtain 24. An action in this task consists of simply applying a mathematical operation to combine two numbers e.g. $2+3=5$, the resulting state from the operation is the set of remaining numbers e.g “5 4 5”.
F.2 Prompts
The prompt used to generate actions for the Game-of-24 task is presented in Figure 15. Likewise, the prompt used to evaluate states is presented in Figure 16. Like WebShop, this evaluation prompt is only used to prompt base models; STL value models are only prompted with the current trajectory. However, unlike WebShop, values are not real numbers between 1 and 10, but rather 0.001, 1, and 20, corresponding to the labels of impossible, likely, and sure that the remaining numbers can be combined to reach 24. Note that these values were used by the original Tree-of-Thoughts paper [64], but are ad-hoc and are used purely as labels. For all value estimates (base model or fine-tuned), we prompt the value model 3 times and use the median score as the state value estimate. During the data generation phase, since we need a single rationale to fine-tune on which to construct the action-outcome rationale, we choose the rationale corresponding to the median of the 3 scores.
F.3 Evaluation
For all tested BFS methods, we use the same setup as the Tree-of-Thoughts paper, i.e., we select 5 of the best actions (set the branching factor to 5) at each of the 4 steps (two numbers are combined during each step).
F.4 Implementing STL
Since the state space is quite limited, we combine training examples from the previous k1k-1 iterations with current examples to train the value model in the $k$th iteration. If the same state is encountered multiple times in different iterations, we defer to the value judgment from the latest iteration.
Unlike with WebShop and HotpotQA, we do not train value models at each depth, due to the small state space. Additionally, we also use a branching factor of 5 during data generation.
Table 6: Performance comparison across different numbers of tasks seen during self-improvement.
Tasks Seen During Improvement Accuracy (25 tasks / iter) Accuracy (50 tasks / iter)
0 38.0 38.0
25 34.0 -
50 30.0 48.0
F.5 Investigating Improvement Dynamics
In §4.3, we claimed that the initial decrease in performance on unseen tasks in Game-of-24 was due to a lack of sufficient tasks seen during self-improvement. We confirmed this claim by rerunning the experiment by rolling out $50$ instead of $25$ tasks per iteration. The results are summarized in Table 6, from which we see that performance actually improves in the first iteration if the number of tasks seen during self-improvement is increased to $50$.
Appendix G Costs
Here we detail how we computed costs in the efficiency analysis in §5.
G.1 Data Generation Costs
To compute data generation costs during STL, we use OpenAI’s 555openai.com/api/pricing for gpt-3.5-turbo and Groq’s 666groq.com/pricing pricing tables for closed source llama models. Note that we only use Groq for a pricing estimate, as inference and training were run on in-house GPUs. As costs may change over time, we provide the pricing figures that we used for our experiments in Table 7. We also note that experiments were initially run on gpt-3.5-turbo-0613 on OpenAI Azure, which is multiple times as expensive as the newer gpt-3.5-turbo-0125 due to resource allocation on Azure, despite being the better model. We therefore choose to report the data generation costs using the newer pricing point as it is more comparable to the current prices of other models.
G.2 Fine-tuning Costs
We account for the costs incurred from fine-tuning llama-3.1-8b-instruct during STL. Fine-tuning with a single A40 GPU takes 4.5 hours for the WebShop task. While it is difficult to estimate costs on our own in-house GPUs, we can estimate the cost by using VastAI’s figure of \$0.39 per A40 GPU hour  [^7] https://vast.ai/pricing/gpu/A40 when these experiments were run in April, 2025. Thus, the total training run cost roughly \$1.76. We also note that as the number of tasks seen at test time scales, the fine-tuning costs become increasingly negligible.
G.3 Inference Costs
To compute inference cost, we use the pricing figures in Table 7.
Table 7: Inference pricing used for cost analysis.
Prompt Tokens (\$ / 1000 tokens) Completion Tokens (\$ / 1000 tokens)
gpt-3.5-turbo 0.0005 0.0015
gpt-4o 0.0025 0.01
llama-3.1-8b-instruct 0.00005 0.00008
Appendix H Model Fine-tuning and Serving
Table 8: Hyperparameters during STL training.
warmup-steps learning-rate weight-decay per-device-batch size lora-r lora-alpha
STL $10$ $2e^{-4}$ $0.01$ 88 $16$ $16$
Table 9: Effect of $\gamma$ on Performance.
$\gamma$ $0.75$ $0.80$ $0.85$ $0.90$ $0.95$ $0.99$ $1.0$
WebShop Success Rate $32.0$ $24.0$ $24.0$ $24.0$ $28.0$ $42.0$ $46.0$
Fine-tuning the value model for STL is carried out on a single NVIDIA A40 GPU. We use LoRA finetuning [25] and use models provided by unsloth888unsloth.ai. The hyperparameters used are in Table 8. We fine-tuned Game-of-24 value models for 10 epochs and WebShop value models for 20 epochs due to the differences in difficulty for models to learn the format of the action and state representations. We find that we require this large number of epochs to learn both the rationale structure and a good representation of transition dynamics. We note that it is clear that overfitting is now happening since unseen task performance does improve across tasks.
Additionally, we serve base and fine-tuned models using vLLM 999docs.vllm.ai/en/latest/ [33] for efficient value estimation of new states during search. We use $\texttt{temperature}=1.0$ and $\texttt{max\_tokens}=3192$.
As mentioned in §3, we use $\gamma=1.0$ based on prior work. We also ran a small hyperparameter tuning validation for $\gamma\in\{0.95,1.0\}$ on a 50-example held-out validation set and found $\gamma=1.0$ had an average reward of 73.5 compared to 72.7 for $\gamma=0.95$.
Finally, we perform a small analysis on the effect of the discount factor $\gamma$ on search performance on the test. The results of the analysis with a gpt-3.5-turbo policy on WebShop are presented in Table 9. The affect of $\gamma$ on the text set is more stark.
Appendix I Significance Testing
In §4.1 and §4.2, we use the paired bootstrap test to test the statistical significance of our experimental results. Following [3], we set $b=10^{6}$. For WebShop, we run the significance test twice: once for score (average reward) and a separate time for success rate.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and the introduction, we claim that using STL to improve a smaller LLM value model can enable search performance matching that of closed-source models, which is experimentally confirmed in the domains of web agent tasks (§4.1), multi-step question answering (§4.2), and math puzzles (§4.3). The abstract and introduction also make the claim that STL enables more efficient search, which is supported in §5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations for our work are presented in Appendix B.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include any theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Yes, our paper fully discloses the information needed to reproduce results. §3 and Algorithm 1 outline the STL approach in detail, while Appendices DE, and F provide the prompts and implementation details necessary for each domain evaluated. Furthermore, we provide cost calculation specifications in Appendix G and fine-tuning hyperparameters in Appendix H. We also provide our code in the supplementary materials.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We have provided a link to the code to reproduce our experiments in §4.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Yes, our paper fully discloses the experimental setting for each domain and evaluated method in §4. This section also specifies the type of search used, the number of tasks used for improvement, and the number iteration of self-improvement used. We further provide prompts and other details for each domain in Appendices DE, and F. Finally, we provide cost calculation specifications in Appendix G and fine-tuning hyperparameters in Appendix H.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We measure statistical significance throughout §4 such as in Table 1 and Table 3 using the paired bootstrap test [3].
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We present compute requirements used for our experiments in Appendices G and H.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, our research does conform to the NeurIPS Code of Ethics. We note that we do not propose a new dataset nor employ human participants.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Both the potential positive and negative societal impacts are discussed in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: We outline safeguards for code release in Appendix A.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We have cited the creators of the original assets, including datasets and models that we use for evaluation in §4.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release any new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: Our work does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development of this work did not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.